### v1

# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

Abstract
We present the first finite-sample analysis for policy evaluation in robust average-reward Markov Decision Processes (MDPs). Prior works in this setting have established only asymptotic convergence guarantees, leaving open the question of sample complexity. In this work, we address this gap by establishing that the robust Bellman operator is a contraction under the span semi-norm, and developing a stochastic approximation framework with controlled bias. Our approach builds upon Multi-Level Monte Carlo (MLMC) techniques to estimate the robust Bellman operator efficiently. To overcome the infinite expected sample complexity inherent in standard MLMC, we introduce a truncation mechanism based on a geometric distribution, ensuring a finite constant sample complexity while maintaining a small bias that decays exponentially with the truncation level. Our method achieves the order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for robust policy evaluation and robust average reward estimation, marking a significant advancement in robust reinforcement learning theory.
1 Introduction
Reinforcement learning (RL) has demonstrated significant success across various domains, including robotics, finance, and healthcare, by enabling agents to learn optimal decision-making strategies through interaction with the environment. However, in many real-world applications, direct interaction with the environment is impractical due to safety concerns, high costs, or data collection constraints (Sünderhauf et al., 2018; Höfer et al., 2021). This challenge is particularly evident in scenarios where agents are trained in simulated environments before being deployed in the real world, such as in robotic control and autonomous driving. The discrepancy between the simulated and real-world environments, known as the simulation-to-reality gap, often leads to performance degradation when the learned policy encounters unmodeled uncertainties. Robust reinforcement learning (robust RL) addresses this challenge by formulating the learning problem as an optimization over an uncertainty set of transition probabilities, ensuring reliable performance under worst-case conditions. In this work, we focus on the problem of evaluating the robust value function and robust average reward for a given policy using only data sampled from a simulator (nominal model), aiming to enhance generalization and mitigate the impact of transition uncertainty in real-world deployment.
Reinforcement learning problems under infinite time horizons are typically studied under two primary reward formulations: the discounted reward setting, where future rewards are exponentially discounted, and the average reward setting, which focuses on optimizing long-term performance. While the discounted-reward formulation is widely used, it may lead to myopic policies that underperform in applications requiring sustained long-term efficiency, such as queueing systems, inventory management, and network control. In contrast, the average-reward setting is more suitable for environments where decisions impact long-term operational efficiency. Despite its advantages, robust reinforcement learning under the average-reward criterion remains largely unexplored. Existing works on robust average-reward RL primarily provide asymptotic guarantees (Wang et al., 2023a, b, 2024), lacking scalable algorithms with finite-time performance bounds. This gap highlights the need for principled approaches that ensure robustness against model uncertainties while maintaining strong long-term performance guarantees.
Solving the robust average-reward reinforcement learning problem is significantly more challenging than its non-robust counterpart, with the primary difficulty arising in policy evaluation. Specifically, the goal is to compute the worst-case value function and worst-case average reward over an entire uncertainty set of transition models while having access only to samples from a nominal transition model. In this paper, we investigate three types of uncertainty sets: Contamination uncertainty sets, TV distance uncertainty sets, and Wasserstein distance uncertainty sets. Unlike the standard average-reward setting, where value functions and average rewards can be estimated directly from observed trajectories, the robust setting introduces an additional layer of complexity due to the need to optimize against adversarial transitions. Consequently, conventional approaches based on direct estimation such as (Wei et al., 2020; Agarwal et al., 2021) immediately fail, as they do not account for the worst-case nature of the problem. Overcoming this challenge requires new algorithmic techniques that can infer the worst-case dynamics using only limited samples from the nominal model.
1.1 Challenges and Contributions
A common approach to policy evaluation in robust RL is to solve the corresponding robust Bellman operator. However, robust average-reward RL presents additional difficulties compared to the robust discounted setting. In the discounted case, the presence of a discount factor induces a contraction property in the robust Bellman operator (Wang and Zou, 2022; Zhou et al., 2024), facilitating stable iterative updates. In contrast, the average-reward Bellman operator lacks a contraction property with respect to any norm even in the non-robust setting (Zhang et al., 2021), making standard fixed-point analysis inapplicable. Due to this fundamental limitation, existing works on robust average-reward RL such as (Wang et al., 2023b) rely on asymptotic techniques, primarily leveraging ordinary differential equation (ODE) analysis to examine the behavior of temporal difference (TD) learning. These methods exploit the asymptotic stability of the corresponding ODE (Borkar, 2023) to establish almost sure convergence but fail to provide finite-sample performance guarantees. Addressing this limitation requires novel analytical tools and algorithmic techniques capable of providing explicit finite-sample bounds for robust policy evaluation and optimization.
In this work, we first establish and exploit a key structural property of the robust average-reward Bellman operator: its contraction under the span semi-norm, denoted as $\|\cdot\|_{\mathrm{sp}}$. This fundamental result enables the use of stochastic approximation techniques similar to (Zhang et al., 2021) to analyze and bound the error in policy evaluation, overcoming the lack of a standard contraction property that has hindered prior finite-sample analyses. Building on this insight, we develop a novel stochastic approximation framework tailored to the robust average-reward setting. Our approach simultaneously estimates both the robust value function and the robust average reward, leading to an efficient iterative procedure for solving the robust Bellman equation. A critical challenge in this framework uner TV and Wasserstein distance uncertainty sets is accurately estimating the worst-case transition effects, which requires computing the support function of the uncertainty set. While previous work has leveraged Multi-Level Monte Carlo (MLMC) for this task, existing MLMC-based estimators suffer from infinite expected sample complexity due to the unbounded nature of the required geometric sampling. To address this, we introduce a truncation mechanism based on a truncated geometric distribution, ensuring that the sample complexity remains finite while maintaining an exponentially decaying bias. With these techniques, we derive the first finite-sample complexity guarantee for policy evaluation in robust average-reward RL, achieving an optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity bound. The main contributions of this paper are summarized as follows:
  • We prove that under the ergodicity assumption, the robust average-reward Bellman operator is a contraction with respect to the span semi-norm (Theorem 2). This key result enables the application of stochastic approximation techniques for policy evaluation.
  • We prove the convergence of stochastic approximation under the span semi-norm contraction and under i.i.d. with noise with non-zero bias (Theorem 3).
  • We develop an efficient method for computing estimates for the robust Bellman operator across TV distance and Wasserstein distance uncertainty sets. By modifying MLMC with a truncated geometric sampling scheme, we ensure finite expected sample complexity while keeping variance controlled and bias decaying exponentially with truncation level (Theorem 4-6).
  • We propose a novel temporal difference learning method that iteratively updates the robust value function and the robust average reward, facilitating efficient policy evaluation in robust average-reward RL. We establish the first non-asymptotic sample complexity result for policy evaluation in robust average-reward RL, proving an order-optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity for policy evaluation (Theorem 7), along with a $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity for robust average reward estimation (Theorem 8).
2 Related Work
The theoretical gauarantees of robust average reward have been studied by the following works. (Wang et al., 2023a) takes a model-based perspective, approximating robust average-reward MDPs with discounted MDPs and proving uniform convergence of the robust discounted value function as the discount factor approaches one, employing dynamic programming and Blackwell optimality arguments to characterize optimal policies. (Wang et al., 2023b) propose a model-free approach by developing robust relative value iteration (RVI) TD and Q-learning algorithms, proving their almost sure convergence using stochastic approximation, martingale theory, and multi-level Monte Carlo estimators to handle non-linearity in the robust Bellman operator. While these studies provide fundamental insights into robust average-reward RL, they do not establish explicit convergence rate guarantees due to the lack of contraction properties in the robust Bellman operator. In addition, (Sun et al., 2024) studies the policy optimization of average reward robust MDPs with assuming direct queries of the value functions.
Policy evaluation in robust discounted-reward reinforcement learning with finite sample gaurauntees has been extensively studied, with recent works (Wang and Zou, 2022; Zhou et al., 2024; Li et al., 2022; Kumar et al., 2023; Kuang et al., 2022) all focusing on solving the robust Bellman equation by finding its fixed-point solution. This approach is made feasible by the contraction property of the robust Bellman operator under the sup-norm, which arises due to the presence of a discount factor $\gamma<1$. However, this fundamental approach does not directly extend to the robust average-reward setting, where the absence of a discount factor removes the contraction property under any norm. As a result, existing robust discounted methods cannot be applied in the robust average reward RL setting.
3 Formulation
3.1 Average-reward MDPs.
[…]
A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps a state to a distribution over $\mathcal{A}$, following which the agent takes action $a$ at state $s$ with probability $\pi(a|s)$. Under a transition kernel $\mathsf{P}$, the average-reward of $\pi$ starting from $s\in\mathcal{S}$ is defined as
$g_{\mathsf{P}}^{\pi}(s)\triangleq\lim_{T\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}\frac{1}{T}\sum^{T-1}_{n=0}r_{t}|S_{0}=s\bigg{]}.$ (1)
The relative value function is defined to measure the cumulative difference between the reward and $g^{\pi}_{\mathsf{P}}$:
$V^{\pi}_{\mathsf{P}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}\sum^{\infty}_{t=0}(r_{t}-g^{\pi}_{\mathsf{P}})|S_{0}=s\bigg{]}.$ (2)
Then $(g^{\pi}_{\mathsf{P}},V^{\pi}_{\mathsf{P}})$ satisfies the following Bellman equation (Puterman, 1994):
$V^{\pi}_{\mathsf{P}}(s)=\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}r(s,A)-g^{\pi}_{\mathsf{P}}(s)+\sum_{s^{\prime}\in\mathcal{S}}p^{A}_{s,s^{\prime}}V^{\pi}_{\mathsf{P}}(s^{\prime})\bigg{]}.$ (3)
3.2 Robust average-reward MDPs.
For robust MDPs, the transition kernel is assumed to be in some uncertainty set $\mathcal{P}$. At each time step, the environment transits to the next state according to an arbitrary transition kernel $\mathsf{P}\in\mathcal{P}$. In this paper, we focus on the $(s,a)$-rectangular compact uncertainty set (Nilim and Ghaoui, 2003; Iyengar, 2005), i.e., $\mathcal{P}=\bigotimes_{s,a}\mathcal{P}^{a}_{s}$, where $\mathcal{P}^{a}_{s}\subseteq\Delta(\mathcal{S})$. Popular uncertainty sets include those defined by the contamination model (Huber, 1965; Wang and Zou, 2022), total variation (Lim et al., 2013), Chi-squared divergence (Iyengar, 2005) and Wasserstein distance (Gao and Kleywegt, 2022).
We investigate the worst-case average-reward over the uncertainty set of MDPs. Specifically, define the robust average-reward of a policy $\pi$ as
$g^{\pi}_{\mathcal{P}}(s)\triangleq\min_{\kappa\in\bigotimes_{n\geq 0}\mathcal{P}}\lim_{T\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{T}\sum^{T-1}_{t=0}r_{t}|S_{0}=s\right],$ (4)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{n\geq 0}\mathcal{P}$. It was shown in (Wang et al., 2023a) that the worst case under the time-varying model is equivalent to the one under the stationary model:
$g^{\pi}_{\mathcal{P}}(s)=\min_{\mathsf{P}\in\mathcal{P}}\lim_{T\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\left[\frac{1}{T}\sum^{T-1}_{t=0}r_{t}|S_{0}=s\right].$ (5)
Therefore, we limit our focus to the stationary model. We refer to the minimizers of (5) as the worst-case transition kernels for the policy $\pi$, and denote the set of all possible worst-case transition kernels by $\Omega^{\pi}_{g}$, i.e., $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$.
We focus on the model-free setting, where only samples from the nominal MDP denoted as $\mathsf{P}$ (the centroid of the uncertainty set) are available. We investigate the problem of robust policy evaluation and robust average reward estimation, which means for a given policy $\pi$, we aim to estimate the robust value function and the robust average reward. We now formally define the robust value function $V^{\pi}_{\mathsf{P}_{V}}$ by connecting it with the following robust Bellman equation:
Theorem 1 (Robust Bellman Equation, Theorem 3.1 in (Wang et al., 2023b)).
If $(g,V)$ is a solution to the robust Bellman equation
$V(s)=\sum_{a}\pi(a|s)\big{(}r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{)},\quad\forall s\in\mathcal{S},$ (6)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)=\min_{p\in\mathcal{P}^{a}_{s}}pV$, then the following properties hold:
  1. 1.
    The scalar $g$ corresponds to the robust average reward, i.e., $g=g^{\pi}_{\mathcal{P}}$.
  2. 2.
    The worst-case transition kernel $\mathsf{P}_{V}$ belongs to the set of minimizing transition kernels, i.e., $\mathsf{P}_{V}\in\Omega^{\pi}_{g}$, where
    $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}.$ (7)
  3. 3.
    The function $V$ is unique up to an additive constant, where if $V$ is a solution to the bellman equation, then we have
    $V=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e},$ (8)
    where $c\in\mathbb{R}$ and $\mathbf{e}$ is the all-ones vector in $\mathbb{R}^{|\mathcal{S}|}$.
This robust Bellman equation characterizes the worst-case value function under the uncertainty set. In particular, $\sigma_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$. Unlike the robust discounted case, where the contraction property of the Bellman operator under the sup-norm enables straightforward fixed-point iteration, the robust average-reward Bellman equation does not induce contraction under any norm, making direct iterative methods inapplicable. We now characterize the explicit forms of $\sigma_{\mathcal{P}^{a}_{s}}(V)$ for different compact uncertainty sets are as follows:
Contamination Uncertainty Set The $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\mathsf{P}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Under this uncertainty set, the support function can be computed as
$\sigma_{\mathcal{P}^{a}_{s}}(V)=(1-\delta)\mathsf{P}^{a}_{s}V+\delta\min_{s}V(s),$ (9)
and this is linear in the nominal transition kernel $\mathsf{P}^{a}_{s}$.
Total Variation Uncertainty Set. The total variation uncertainty set is $\mathcal{P}^{a}_{s}=\{q\in\Delta(|\mathcal{S}|):\frac{1}{2}\|q-\mathsf{P}^{a}_{s}\|_{1}\leq\delta\},$ define $\|\cdot\|_{\mathrm{sp}}$ as the span semi-norm and the support function can be computed using its dual function Iyengar (2005):
$\sigma_{\mathcal{P}^{a}_{s}}(V)$ $=\max_{\mu\geq\mathbf{0}}\big{(}\mathsf{P}^{a}_{s}(V-\mu)-\delta\|V-\mu\|_{\mathrm{sp}}\big{)}.$ (10)
Wasserstein Distance Uncertainty Sets. Consider the metric space $(\mathcal{S},d)$ by defining some distance metric $d$. For some parameter $l\in[1,\infty)$ and two distributions $p,q\in\Delta(\mathcal{S})$, define the $l$-Wasserstein distance between them as $W_{l}(q,p)=\inf_{\mu\in\Gamma(p,q)}\|d\|_{\mu,l}$, where $\Gamma(p,q)$ denotes the distributions over $\mathcal{S}\times\mathcal{S}$ with marginal distributions $p,q$, and $\|d\|_{\mu,l}=\big{(}\mathbb{E}_{(X,Y)\sim\mu}\big{[}d(X,Y)^{l}\big{]}\big{)}^{1/l}$. The Wasserstein distance uncertainty set is then defined as
$\mathcal{P}^{a}_{s}=\left\{q\in\Delta(|\mathcal{S}|):W_{l}(\mathsf{P}^{a}_{s},q)\leq\delta\right\}.$ (11)
The support function w.r.t. the Wasserstein distance set, can be calculated as follows (Gao and Kleywegt, 2023):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\sup_{\lambda\geq 0}\left(-\lambda\delta^{l}+\mathbb{E}_{\mathsf{P}^{a}_{s}}\big{[}\inf_{y}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}\big{]}\right).$ (12)
4 Robust Bellman Operator
Motivated by Theorem 1, we define the robust Bellman operator, which forms the basis for our policy evaluation procedure.
Definition 1 (Robust Bellman Operator, Wang et al. (2023b)).
The robust Bellman operator $\mathbf{T}_{g}$ is defined as:
$\mathbf{T}_{g}(V)(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{]},\quad\forall s\in\mathcal{S}.$ (13)
The operator $\mathbf{T}_{g}$ transforms a candidate value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under the span semi-norm. This allows us to develop provably efficient stochastic approximation algorithms. Throughout this paper, we make the following standard assumption regarding the structure of the induced Markov chain.
Assumption 1.
The Markov chain induced by $\pi$ is irreducible and aperiodic for all $\mathsf{P}\in\mathcal{P}$.
Assumption 1 is used widely in all robust average reinforcement learning literatures (Wang et al., 2023a, b, 2024; Sun et al., 2024). This assumption ensures that, under any transition model within the uncertainty set, the policy $\pi$ induces a single recurrent communicating class. A well-known result in average-reward MDPs states that under Assumption 1, the average reward is independent of the starting state, i.e., for any $\mathsf{P}\in\mathcal{P}$ and all $s,s^{\prime}\in\mathcal{S}$,
$g^{\pi}_{\mathsf{P}}(s)=g^{\pi}_{\mathsf{P}}(s^{\prime}).$ (14)
Thus, we can drop the dependence on the initial state and simply write $g^{\pi}_{\mathsf{P}}$ as the robust average reward.
Under Assumption 1, we are able to establish the semi-norm contraction property. Before proceeding, we first establish the semi-norm property of non-robust average reward bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows.
$\mathbf{T}_{g}^{\mathsf{P}}(V)(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g+\sum_{s^{\prime}}\mathsf{P}(s^{\prime}|s,a)V(s^{\prime})\big{]},\quad\forall s\in\mathcal{S}.$ (15)
Lemma 1.
Let $\mathcal{S}$ be a finite state space, and let $\pi$ be a stationary policy. If the Markov chain induced by $\pi$ under the transition $\mathsf{P}$ is irreducible and aperiodic, there exists a constant $\beta\in(0,1)$ such that for all $V_{1},V_{2}\in\mathbb{R}^{S}$ and $g\in\mathbb{R}$
$\|\mathbf{T}_{g}^{\mathsf{P}}(V_{1})-\mathbf{T}_{g}^{\mathsf{P}}(V_{2})\|_{\mathrm{sp}}\leq\beta\|V_{1}-V_{2}\|_{\mathrm{sp}},$ (16)
where
$\|v\|_{\mathrm{sp}}\coloneqq\max_{s}\,v(s)-\min_{s}\,v(s).$
The proof of Lemma 1 is in Appendix A.1, where the properties of irreduible and aperiodic finite state Markov chain is utilized. Thus we show the (non-robust) average reward bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under the span semi-norm. Based on the above results, we now formally establish the contraction property of the robust average reward bellman operator by leveraging Lemma 1 and the compactness of the uncertainty sets.
Theorem 2.
Under Assumption 1, and if $\mathcal{P}$ is compact, the robust bellman operator $\mathbf{T}_{g}$ is a contractive mapping with respect to the span semi-norm for any $g$. Specifically, there exist $\gamma\in(0,1)$ such that
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathrm{sp}}\leq\gamma\|V_{1}-V_{2}\|_{\mathrm{sp}},\quad\forall V_{1},V_{2}\in\mathbb{R}^{S},g\in\mathbb{R}$ (17)
where
$\|v\|_{\mathrm{sp}}\coloneqq\max_{s}\,v(s)-\min_{s}\,v(s).$
The proof of Theorem 2 is in Appendix A.2. Since all the uncertainty sets listed in Section 3.2 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the comtraction property in Theorem 2.
5 Convergence of Span Contraction with Bias
In the previous section, we established that the robust Bellman operator is a contraction under the span semi-norm, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 6). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which is further used to establish the sample complexity of policy evaluation in robust average reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
$x^{t+1}=x^{t}+\eta_{t}\bigl{[}\widehat{H}(x^{t})-x^{t}\bigr{]},\quad\text{where}\quad\widehat{H}(x^{t})=H(x^{t})+w^{t}.$ (18)
with $\eta_{t}>0$ being the step-size sequance and with the following assumptions on the operator $H$ and noise $\omega^{t}$:
  • $H$ is a contractive mapping with respect to the span semi-norm, there exist $\gamma\in(0,1)$ such that
    $\|H(x)-H(y)\|_{\mathrm{sp}}\leq\gamma\,\|x-y\|_{\mathrm{sp}},\quad\forall x,y$ (19)
  • the noise terms $\omega^{t}$ are i.i.d. and have bounded bias and variance
    $\mathbb{E}[\,\|w^{t}\|_{\mathrm{sp}}^{2}|\mathcal{F}^{t}]\leq A+B\,\|x^{t}-x^{*}\|_{\mathrm{sp}}^{2}\quad\text{and}\quad\bigl{\|}\mathbb{E}[\,w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathrm{sp}}\leq\varepsilon_{\mathrm{bias}}$ (20)
Theorem 3.
If $x^{t}$ is generated by (18) with all assumptions in (19) and (20) satisfied, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}\right)\|x^{0}-x^{*}\|^{2}_{\mathrm{sp}}+\operatorname*{\mathcal{O}}\left(\frac{A}{(1-\gamma)^{2}T}\right)+\operatorname*{\mathcal{O}}\left(\frac{x_{\mathrm{sp}}\varepsilon_{\text{bias}}\log T}{1-\gamma}\right)$ (21)
where $x_{\mathrm{sp}}\coloneqq\sup_{x}\|x\|_{\mathrm{sp}}$ is the upper bound of the span for all $x^{t}$.
Theorem 3 adapts the analysis of (Zhang et al., 2021) and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms (see (59)-(66) in Appendix B.3) to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$. The detailed proof of Theorem 3 along with the exact constant terms is in Appendix B.
6 Queries from Uncertainty Set
In this section, we aim to construct an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ in various uncertainty sets. Recall that the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$ as defined in the robust Bellman equation in Theorem 1. The explicit forms of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for different uncertainty sets were characterized in (9)-(12). Our goal in this section is to construct efficient estimators $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ that approximates ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ while maintaining controlled variance and finite sample complexity.
6.1 Linear Contamination Uncertainty Set
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\mathsf{P}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Since the support function can be computed by (9) and the expression is linear in the nominal transition kernel $\mathsf{P}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
$\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\triangleq(1-\delta)V(s^{\prime})+\delta\min_{x}V(x),$ (22)
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (22) is just one. A well know result from (Wang et al., 2023b) is that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (22) is unbiased and has bounded variance as follows:
$\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]={\sigma}_{\mathcal{P}^{a}_{s}}(V),\quad\text{and}\quad\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\|V\|^{2}$ (23)
6.2 Non-Linear Uncertainty Sets
Non-linear uncertainty sets such as TV distance uncertainty set and Wasserstein distance uncertainty sets have a non-linear relationship between the nonminal distribution $\mathsf{P}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (Blanchet and Glynn, 2015; Blanchet et al., 2019; Wang et al., 2023b) have proposed a multi-level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approach all require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator:
$\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$ (24)
To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a level-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
Algorithm 1 Truncated MLMC Estimator for TV and Wasserstein Unceretainty Sets
Input: $s\in\mathcal{S}$, $a\in\mathcal{A}$, Truncation level $N_{\max}$, Value function $V$
1:Sample $N\sim\mathrm{Geom}(0.5)$
2:$N^{\prime}\leftarrow\min\{N,N_{\max}\}$
3:Collect $2^{N^{\prime}+1}$ i.i.d. samples of $\{s^{\prime}_{i}\}^{2^{N^{\prime}+1}}_{i=1}$ with $s^{\prime}_{i}\sim\mathsf{P}^{a}_{s}$ for each $i$
4:$\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i}\}}$
5:$\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i-1}\}}$
6:$\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}+1}}\sum_{i=1}^{2^{N^{\prime}+1}}\mathbbm{1}_{\{s^{\prime}_{i}\}}$
7:$\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}\leftarrow\mathbbm{1}_{\{s^{\prime}_{1}\}}$
8:if TV distance uncertainty set then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (10)
9:else if Wasserstein distance uncertainty set then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (12)
10:$\Delta_{N^{\prime}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V)-\frac{1}{2}\Bigl{[}\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V)+\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)\Bigr{]}$
11:$\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)},\text{where }p^{\prime}(n)=\mathbb{P}(N^{\prime}=n)$ return $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$
In particular, if $n<N_{\max}$, then $\{N^{\prime}=n\}=\{N=n\}$ with probability $(\tfrac{1}{2})^{n+1}$, while $\{N^{\prime}=N_{\max}\}$ has probability $\sum_{m=N_{\max}}^{\infty}(1/2)^{m+1}=2^{-N_{\max}}$. After obtaining $N^{\prime}$, Algorithm 1 then collects a set of $2^{N^{\prime}+1}$ i.i.d. samples from the nominal transition model to construct empirical estimators for different transition distributions. The core of the approach lies in computing the support function estimates for TV and Wasserstein uncertainty sets using a correction term $\Delta_{N^{\prime}}(V)$, which accounts for the bias introduced by truncation. This correction ensures that the final estimator maintains a low bias while achieving a finite sample complexity. We now present several crucial properties of Algorithm 1.
6.2.1 Sample Complexity for Querying Non-Linear Uncertainty Sets
Theorem 4 (Finite Sample Complexity).
Under Algorithm 1, denote $M=2^{N^{\prime}+1}$ as the random number of samples (where $N^{\prime}=\min\{N,N_{\max}\}$). Then
$\mathbb{E}[M]=N_{\max}+2=\mathcal{O}(N_{\max}).$ (25)
The proof of Theorem 4 is in Appendix C.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
6.2.2 Exponential Bias Decay
Theorem 5 (Exponentially Decaying Bias).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 the under TV uncertainty set, we have:
$\left|\mathbb{E}\bigl{[}\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\bigr{]}\right|\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (26)
and under Wasserstein uncertainty set, we have:
$\left|\mathbb{E}\bigl{[}\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\bigr{]}\right|\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (27)
Theorem 5 establishes that the bias of the truncated MLMC estimator decays exponentially with $N_{\max}$, ensuring that truncation does not significantly affect accuracy. This result follows from observing that the deviation introduced by truncation can be expressed as a sum of differences between support function estimates at different level, and each of which is controlled by the $\ell_{1}$-distance between transition distributions. Thus, we can use binomial concentration property to ensure the exponentially decaying bias.
The proof of Theorem 5 is in Appendix C.2. One important lemma used in the proof is the following Lemma 2, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
Lemma 2.
For any $p,q\in\Delta(\mathcal{S})$, let $\mathcal{P}_{TV}$ and $\mathcal{Q}_{TV}$ denote the TV distance uncertainty set with radius $\delta$ centering at $p$ and $q$ respectively, and let $\mathcal{P}_{W}$ and $\mathcal{Q}_{W}$ denote the Wasserstein distance uncertainty set with radius $\delta$ centering at $p$ and $q$ respectively. Then for any value function $V$, we have:
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (28)
$|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (29)
We refer the proof of Theorem 5 to Appendix C.3.
6.2.3 Linear Variance
Theorem 6 (Linear Variance).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 then under TV distance uncertainty set, we have:
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq 3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (30)
and under Wasserstein distance uncertainty set, we have:
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq 3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (31)
Theorem 6 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 6 is in Appendix C.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
7 Propoosed Algorithm and Final Results
We present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section 5, ensuring both efficient sample complexity and finite-time convergence guarantees.
The algorithm is divided into two main phases. The first phase (Lines 1-7) estimates the robust value function. The noisy Bellman operator is computed using the estimator $\hat{\sigma}_{\mathcal{P}_{s}^{a}}(V_{t})$ obtained depending on the uncertainty set type. Then the iterative update follows a stochastic approximation scheme with step size $\eta_{t}$, ensuring convergence while maintaining stability. Finally, the value function is centered at an anchor state $s_{0}$ to remove the ambiguity due to its additive invariance. The second phase (Lines 8-14) estimates the robust average reward by utilizing $V_{T}$ from the output of the first phase. The expected Bellman residual $\delta_{t}(s)$ is computed across all states and averaging it to obtain $\bar{\delta}_{t}$. A separate stochastic approximation update with step size $\beta_{t}$ is then applied to refine $g_{t}$, ensuring convergence to the robust worst-case average reward. By combining these two phases, Algorithm 2 provides an efficient and provably convergent method for robust policy evaluation under average-reward criteria, marking a significant advancement over prior methods that only provided asymptotic guarantees.
Algorithm 2 Robust Average Reward TD
Input: Policy $\pi$, Initial values $V_{0}$, $g_{0}=0$, Stepsizes $\eta_{t}$, $\beta_{t}$, Truncation level $N_{\max}$, $t=0,1,\ldots,T-1$, Anchor state $s_{0}\in\mathcal{S}$
1:for $t=0,1,\ldots,T-1$ do
2:     for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
3:         if Contamination uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (22)
4:         else if TV distance or Wasserstein distance uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
5:     $\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\leftarrow\sum_{a}\pi(a|s)\big{[}r(s,a)-g_{0}+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\big{]},\quad\forall s\in\mathcal{S}$
6:     $V_{t+1}(s)\leftarrow V_{t}(s)+\eta_{t}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-V_{t}(s)\right),\quad\forall s\in\mathcal{S}$
7:     $V_{t+1}(s)=V_{t+1}(s)-V_{t+1}(s_{0}),\quad\forall s\in\mathcal{S}$
8:for $t=0,1,\ldots,T-1$ do
9:     for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
10:         if Contamination uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (22)
11:         else if TV distance or Wasserstein distance uncertainty set then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
12:     $\hat{\delta}_{t}(s)\leftarrow\sum_{a}\pi(a|s)\big{[}r(s,a)+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\big{]}-V_{T}(s),\quad\forall s\in\mathcal{S}$
13:     $\bar{\delta}_{t}\leftarrow\frac{1}{S}\sum_{s}\hat{\delta}_{t}(s)$
14:     $g_{t+1}\leftarrow g_{t}+\beta_{t}(\bar{\delta}_{t}-g_{t})$ return $V_{T}$, $g_{T}$
To derive the sample complexity of robust policy evaluation, we utilize the span semi-norm contraction property of the bellman operator in Theorem 2, and fit Algorithm 2 into the general biased stochastic approximation result in Theorem 3 while incorporating the bias analysis characterized in Section 6. Since each phase of Algorithm 2 contains a loop of length $T$ with all the states and actions updated together, the total samples needed for the entire algorithm in expectation is $2SAT\mathbb{E}[N_{\max}]$, where $\mathbb{E}[N_{\max}]$ is one for contamination uncertainty sets and is $\operatorname*{\mathcal{O}}(N_{\max})$ from Theorem 4 for TV and Wasserstein distance uncertainty sets.
Theorem 7.
If $V_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $V_{T}$.
Theorem 8.
If $g_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\beta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $g_{T}$.
The formal version of Theorem 7 and Theorem 8 along with the proofs are in Appendix D. Theorem 7 and Theorem 8 provide the order-optimal sample complexity of $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ for Algorithm 2 to achieve an $\epsilon$-accurate estimate of $V_{T}$ and $g_{T}$. The proof of Theorem 8 extends the analysis of Theorem 7 to robust average reward estimation. The key difficulty lies in controlling the propagation of error from value function estimates to reward estimation. By again leveraging the contraction property and appropriately tuning step sizes, we establish an $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ complexity bound for robust average reward estimation.
8 Conclusion
This paper provides the first finite-sample analysis for policy evaluation in robust average-reward MDPs, bridging a gap where only asymptotic guarantees existed. By introducing a biased stochastic approximation framework and leveraging the properties of various uncertainty sets, we establish finite-time convergence under biased noise. Our algorithm achieves an order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for policy evaluation, despite the added complexity of robustness.
A crucial step in our analysis is proving that the robust Bellman operator is contractive under the span semi-norm, ensuring the validity of stochastic approximation updates. We further develop a truncated multi-level Monte Carlo estimator that efficiently computes worst-case value functions under total variation and Wasserstein uncertainty, while keeping bias and variance controlled. Our results confirm that robust policy evaluation can be achieved with near-optimal efficiency, comparable to standard stochastic approximation methods.
9 Acknowledgement
We would like to thank Zaiwei Chen for assisting in identifying relevant literature and providing constructive feedback.
Appendix A Span Semi-Norm Contraction Property of Bellman Operator
A.1 Proof of Lemma 1
Let $V_{1},V_{2}:\mathcal{S}\to\mathbb{R}$ and define $\Delta=V_{1}-V_{2}$. Denote $\mathsf{P}^{\pi}$ as the transition matrix of the Markov chain under policy $\pi$. By the definition of the Bellman operator under the transition $\mathsf{P}^{\pi}$, we have
$\mathbf{T}_{g}^{\pi}(V_{1})(s)-\mathbf{T}_{g}^{\pi}(V_{2})(s)=\sum_{s^{\prime}\in\mathcal{S}}\mathsf{P}^{\pi}(s^{\prime}|s)\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}=\mathsf{P}^{\pi}\,\Delta(s).$ (32)
Hence
$\bigl{\|}\mathbf{T}_{g}^{\pi}(V_{1})-\mathbf{T}_{g}^{\pi}(V_{2})\bigr{\|}_{\mathrm{sp}}=\|\mathsf{P}^{\pi}\,\Delta\|_{\mathrm{sp}}.$ (33)
Note that irreducibility and aperiodicity of the Markov chain on a finite state space implies (Norris, 1998) the existence of an integer $m\geq 1$ and $\gamma>0$ such that
$(\mathsf{P}^{\pi})^{m}(s_{1},s_{2})\geq\gamma\quad\forall\,s_{1},s_{2}\in\mathcal{S}.$ (34)
Thus each row of $(\mathsf{P}^{\pi})^{m}$ is a probability distribution with at least positive mass $\gamma$ over all states. We now introduce the following lemma to characterize the contraction property of $(\mathsf{P}^{\pi})^{m}$,
Lemma 3.
Let $M$ be an $S\times S$ stochastic matrix with the property
$M(i,t)\;\geq\;\gamma\quad\text{for all }i,t\in\{1,\dots,S\}\quad\text{and someconstant }0<\gamma\leq 1/S.$
Then for any function $x:\{1,\dots,S\}\to\mathbb{R}$, the span of $Mx$ is strictly reduced:
$\|Mx\|_{\mathrm{sp}}\leq(1-2\,\gamma)\,\|x\|_{\mathrm{sp}},$ (35)
where
$\|x\|_{\mathrm{sp}}=\max_{1\leq s\leq S}\,x(s)-\min_{1\leq s\leq S}\,x(s).$ (36)
Proof.
We first define
$x_{\max}=\max_{s}x(s),\quad x_{\min}=\min_{s}x(s),\quad\|x\|_{\mathrm{sp}}=x_{\max}-x_{\min}.$
If $\|x\|_{\mathrm{sp}}=0$, the statement is trivial since both sides are zero. Otherwise, consider the normalized function
$y(s)\coloneqq\frac{x(s)-x_{\min}}{\,x_{\max}-x_{\min}\,},$
so that $0\leq y(s)\leq 1$ for all $s$ and $\|y\|_{\mathrm{sp}}=1$. Then denote $e$ as the all-ones vectors and we have
$\|Mx\|_{\mathrm{sp}}$ $=\|x_{\min}\cdot Me+(x_{\max}-x_{\min})M\bigl{[}\tfrac{x-x_{\min}}{x_{\max}-x_{\min}}\bigr{]}\|_{\mathrm{sp}}$ (37)
$=(x_{\max}-x_{\min})\bigl{\|}My\bigr{\|}_{\mathrm{sp}}$
$=\|x\|_{\mathrm{sp}}\|My\|_{\mathrm{sp}}.$
It remains to prove $\|My\|_{\mathrm{sp}}\leq(1-2\gamma)$ for $y$ taking the values in $[0,1]$.
Assume $y:\{1,\dots,S\}\to[0,1]$ with $\max_{s}y(s)-\min_{s}y(s)=1$, so in fact $\min_{s}y(s)=0$ and $\max_{s}y(s)=1$. For $1\leq i\leq S$, define
$(My)(i)=\sum_{t=1}^{S}M(i,t)y(t).$ (38)
Let $i^{*}\coloneqq\arg\max_{i}(My)(i)$ and $j^{*}\coloneqq\arg\min_{j}(My)(j)$. Then
$\|My\|_{\mathrm{sp}}=(My)(i^{*})-(My)(j^{*}).$ (39)
To bound this difference, construct a pair of random variables $(T,T^{\prime})$ as follows:
$T\sim M(i^{*},\cdot),\quad T^{\prime}\sim M(j^{*},\cdot),$ (40)
However, we can couple them in such a way that
$\mathbb{P}[T=T^{\prime}]\geq 2\gamma.$
Such a coupling is possible because each row $M(i^{*},t)$ and $M(j^{*},t)$ has at least $\gamma$ mass on every $t$. Under this coupling:
$(My)(i^{*})=\mathbb{E}\bigl{[}y(T)\bigr{]},\quad(My)(j^{*})=\mathbb{E}\bigl{[}y(T^{\prime})\bigr{]}.$
Hence
$(My)(i^{*})-(My)(j^{*})=\mathbb{E}\bigl{[}y(T)-y(T^{\prime})\bigr{]}.$ (41)
But $y(t)\in[0,1]$. Whenever $T=T^{\prime}$, the difference $y(T)-y(T^{\prime})$ is zero. Since $\text{Pr}\left[[\right]T=T^{\prime}]\geq 2\gamma$, the event $T\neq T^{\prime}$ has probability at most $(1-2\gamma)$. On $T\neq T^{\prime}$, the difference $|y(T)-y(T^{\prime})|$ is at most $1$ (because $0\leq y(\cdot)\leq 1$). Therefore
$\bigl{|}(My)(i^{*})-(My)(j^{*})\bigr{|}=\bigl{|}\mathbb{E}\bigl{[}y(T)-y(T^{\prime})\bigr{]}\bigr{|}\leq(1-2\gamma)\cdot 1=1-2\gamma.$ (42)
Thus $\|My\|_{\mathrm{sp}}\leq 1-2\gamma$ whenever $\|y\|_{\mathrm{sp}}=1$ and $y\in[0,1]$. ∎
Directly applying $(\mathsf{P}^{\pi})^{m}$ to Lemma 3 would obtain
$\|(P^{\pi})^{m}\Delta\|_{\mathrm{sp}}\leq(1-2\gamma)\|\Delta\|_{\mathrm{sp}},$ (43)
Where $\gamma$ is defined in (34). Thus there also exist $\beta\in(0,1)$ where $\beta=(1-2\gamma)^{\frac{1}{m}}$ such that $\|(P^{\pi})\Delta\|_{\mathrm{sp}}\leq\beta\|\Delta\|_{\mathrm{sp}}$. Thus, this proves the lemma.
A.2 Proof of Theorem 2
We note that for any $g$, for any $V_{1}$ and $V_{2}$, for all $s\in\mathcal{S}$, we have:
$\mathbf{T}_{g}(V_{1})(s)-\mathbf{T}_{g}(V_{2})(s)$ $=\sum_{a\in\mathcal{A}}\pi(a|s)[\sigma_{p^{a}_{s}}(V_{1})-\sigma_{p^{a}_{s}}(V_{2})]$
$=\sum_{a\in\mathcal{A}}\pi(a|s)[\min_{p\in\mathcal{P}^{a}_{s}}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\min_{p\in\mathcal{P}^{a}_{s}}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\max_{p\in\mathcal{P}^{a}_{s}}\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$ (44)
Where the last inequality comes from the fact that $\min f(x)-\min g(x)\leq\max(f(x)-g(x))$. denote $\tilde{p}(V_{1},V_{2})=\operatornamewithlimits{argmax}_{p\in\mathcal{P}^{a}_{s}}[\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$, we further obtain
$\mathbf{T}_{g}(V_{1})(s)-\mathbf{T}_{g}(V_{2})(s)$ $\leq\sum_{a\in\mathcal{A}}\pi(a|s)\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)[V_{1}(s^{\prime})-V_{2}(s^{\prime})]$ (45)
Since $\tilde{p}(V_{1},V_{2})\in\mathcal{P}$ for all $V_{1},V_{2}$, the above problem transforms to the same as the non-robust setting. Thus, by Lemma 1, there exists $\beta(V_{1},V_{2})\in(0,1)$ such that
${\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|}_{\mathrm{sp}}\leq\left\lVert\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\tilde{p}(s^{\prime}|s,a)[V_{1}(s^{\prime})-V_{2}(s^{\prime})]\right\rVert_{\mathrm{sp}}\leq\beta(V_{1},V_{2}){\|V_{1}-V_{2}\|}_{\mathrm{sp}}$ (46)
Give that $\mathcal{P}$ is compact, the supremum is attainable and there exist $\gamma\in(0,1)$ such that for all $V_{1},V_{2}$, we have ${\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|}_{\mathrm{sp}}\leq\gamma{\|V_{1}-V_{2}\|}_{\mathrm{sp}}$, which concludes the proof.
Appendix B Biased Semi-Norm Stochastic Approximation
We perform analysis of the biased‐noise extension to the span semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift (Zhang et al., 2021) and using properties in dual norms for managing the bias.
B.1 Setup and Notation.
In this section, we override the notation of the span semi-norm by re-writing it as the sup-norm to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalent class $\delta_{\overline{E}}$, define the indicator function $\delta_{\overline{E}}$ as
$\delta_{\bar{E}}(x):=\begin{cases}0&x\in\bar{E},\\ \infty&\text{{otherwise}.}\end{cases}$ (47)
then by (Zhang et al., 2021), the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalent class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
$\|x\|_{c,\overline{E}}\coloneqq(\|\cdot\|_{c}\ast_{\inf}\delta_{\overline{E}})(x)=\inf_{y}(\|x-y\|_{c}+\delta_{\overline{E}}(y))=\inf_{e\in\overline{E}}\|x-e\|_{c}\quad\forall x.$ (48)
Where $\ast_{\inf}$ denotes the infimal convolution operator. Throughout the remaining section, we let $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. Then from (Gupta et al., 2015), we have for any $x$,
$\|x\|_{\mathrm{sp}}=2\|x\|_{\infty,\overline{E}}.$ (49)
We thus restate our problem of analyzing the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$, with the operator $H:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying the contraction property as follows:
$\|H(x)-H(y)\|_{\infty,\overline{E}}\leq\gamma\|x-y\|_{\infty,\overline{E}},\quad\gamma\in(0,1),\quad\forall x,y$ (50)
The stochastic approximation iteration being used is as follows
$x^{t+1}=x^{t}+\eta_{t}\bigl{[}\widehat{H}(x^{t})-x^{t}\bigr{]},\quad\text{where}\quad\widehat{H}(x^{t})=H(x^{t})+w^{t}.$ (51)
We assume:
  • $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$ (In the robust average reward TD case, $B=0$).
  • $\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\infty,\overline{E}}\leq\varepsilon_{\mathrm{bias}}$.
  • $\eta_{t}>0$ is a chosen step‐size sequence (decreasing or constant).
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of (Zhang et al., 2021). Thus, we emphasize mostly on managing the bias.
B.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By (Zhang et al., 2021, Proposition 1–2), using the Moreau envelope function $M(x)$ in Definition 2.2 of (Chen et al., 2020), we define
$M_{\overline{E}}(x)=\bigl{(}M\ast_{\inf}\delta_{\overline{E}}\bigr{)}(x),$
so that there exist $c_{l},c_{u}>0$ with
$c_{l}M_{\overline{E}}(x)\leq\frac{1}{2}\|x\|_{\infty,\overline{E}}^{2}\leq c_{u}M_{\overline{E}}(x),$ (52)
and $M_{\overline{E}}$ is $L$-smooth w.r.t. another semi-norm $\|\cdot\|_{s,\overline{E}}$. Concretely, $L$-smoothness means:
$M_{\overline{E}}(y)\leq M_{\overline{E}}(x)+\langle\nabla M_{\overline{E}}(x),\,y-x\rangle+\tfrac{L}{2}\,\|\,y-x\|_{s,\overline{E}}^{2},\quad\forall\,x,y.$ (53)
Moreover, the gradient of $M_{\overline{E}}$ satisfies $\langle\nabla M_{\overline{E}}(x),\;c\,\mathbf{e}\rangle=0$ for all $x$, and the dual norm denoted as $\|\cdot\|_{*,s,\overline{E}}$ is also L-smooth:
$\|\nabla M_{\overline{E}}(x)-\nabla M_{\overline{E}}(y)\|_{*,s,\overline{E}}\leq L\|\,y-x\|_{s,\overline{E}},\quad\forall\,x,y.$ (54)
Note that since $\|\cdot\|_{s,\overline{E}}$ and $\|\cdot\|_{\infty,\overline{E}}$ are semi-norms on a finite‐dimensional space with the same kernel, there exist $\rho_{1},\rho_{2}>0$ such that
$\rho_{1}\,\|z\|_{\infty,\overline{E}}\leq\|z\|_{s,\overline{E}}\leq\rho_{2}\,\|z\|_{\infty,\overline{E}},\;\forall\,z.$ (55)
Likewise, their dual norms (denoted $\|\cdot\|_{*,s,\overline{E}}$ and $\|\cdot\|_{*,\infty,\overline{E}}$) satisfy the following:
$\frac{1}{\rho_{2}}\|z\|_{*,s,\overline{E}}\leq\|z\|_{*,\infty,\overline{E}}\leq\frac{1}{\rho_{1}}\|z\|_{*,s,\overline{E}},\;\forall\,z.$ (56)
B.3 Formal Statement of Theorem 3
By $L$‐smoothness w.r.t. $\|\cdot\|_{s,\overline{E}}$ in (53), for each $t$,
$M_{\overline{E}}(x^{t+1}-x^{*})\leq M_{\overline{E}}(x^{t}-x^{*})+\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\bigr{\rangle}+\tfrac{L}{2}\|\,x^{t+1}-x^{t}\|_{s,\overline{E}}^{2}.$ (57)
where $x^{t+1}-x^{t}=\eta_{t}[\widehat{H}(x^{t})-x^{t}]=\eta_{t}[H(x^{t})+w^{t}-x^{t}]$. Taking expectation of the second term of the RHS of (57) conditioned on the filtration $\mathcal{F}^{t}$ we obtain,
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]$ $=\eta_{t}\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}+\omega^{t}\rangle|\mathcal{F}^{t}]$
$=\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\omega^{t}\rangle|\mathcal{F}^{t}]$
$=\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[\omega^{t}|\mathcal{F}^{t}]\rangle.$ (58)
[…]
$\bigl{\langle}u,\,v\bigr{\rangle}\leq\|u\|_{*}\,\|v\|,\quad\forall\,u,v.$ (59)
In the biased noise setting, $u=\nabla M_{\overline{E}}(x^{t}-x^{*})$ and $v=\mathbb{E}[w^{t}|\mathcal{F}^{t}]$, with $\|\cdot\|=\|\cdot\|_{\infty,\overline{E}}$. So
$\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq\bigl{\|}\nabla M_{\overline{E}}(x^{t}-x^{*})\bigr{\|}_{*,\,\infty,\overline{E}}\cdot\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\infty,\overline{E}}.$ (60)
Since $\|\mathbb{E}[w^{t}|\mathcal{F}^{t}]\|_{\infty,\overline{E}}\leq\varepsilon_{\text{bias}}$, it remains to bound $\|\nabla M_{\overline{E}}(x^{t}-x^{*})\|_{*,\infty,\overline{E}}$. By setting $y=0$ in (54), we get
$\|\nabla M_{\overline{E}}(x)-\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}\leq L\|x\|_{s,\overline{E}},\quad\forall\,x.$ (61)
Thus,
$\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}\leq\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}},\quad\forall\,x.$ (62)
By (56), we know that there exists $\frac{1}{\rho_{2}}\leq\alpha\leq\frac{1}{\rho_{1}}$ such that
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}$ (63)
Thus, combining (62) and (63) would give:
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}}\big{)},\quad\forall\,x.$ (64)
By (55), we know that $\|x\|_{s,\overline{E}}\leq\|x\|_{\infty,\overline{E}}$, thus we have:
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\rho_{2}\|x\|_{\infty,\overline{E}}\big{)},\quad\forall\,x.$ (65)
Hence, combining the above with (60), there exist some
$G=\mathcal{O}\big{(}\frac{1}{\rho_{1}}\max\{L\rho_{2},\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}\}\big{)}$ (66)
such that
$\mathbb{E}\Bigl{[}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\,w^{t}\rangle|\mathcal{F}^{t}\Bigr{]}=\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq G\bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\bigr{)}\varepsilon_{\text{bias}}.$ (67)
Combining (67) with (B.3) we obtain
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]\leq\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\Bigr{)}$ (68)
To bound the first term in the RHS of (68), note that
$\langle\nabla M_{\bar{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle$ $=\langle\nabla M_{\bar{E}}(x^{t}-x^{*}),H(x^{t})-x^{*}+x^{*}-x^{t}\rangle$
$\overset{(a)}{\leq}M_{\bar{E}}(H(x^{t})-x^{*})-M_{\bar{E}}(x^{t}-x^{*})$
$\overset{(b)}{\leq}\frac{1}{2c_{l}}\|H(x^{t})-H(x^{*})\|^{2}_{c,\bar{E}}-M_{\bar{E}}(x^{t}-x^{*})$
$\overset{(c)}{\leq}\frac{\gamma^{2}}{2c_{l}}\|x^{t}-x^{*}\|^{2}_{c,\bar{E}}-M_{\bar{E}}(x^{t}-x^{*})$
$\leq\left(\frac{\gamma^{2}c_{u}}{c_{l}}-1\right)M_{\bar{E}}(x^{t}-x^{*})$
$\leq-(1-\gamma\sqrt{c_{u}/c_{l}})M_{\bar{E}}(x^{t}-x^{*}),$ (69)
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalent class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (B.3). (68) and Lemma 4 with (57), we arrive as follows:
$\mathbb{E}\Bigl{[}M_{\overline{E}}(x^{t+1}-x^{*})|\mathcal{F}_{t}\Bigr{]}$ $\leq(1-2\alpha_{2}\eta_{t}+\alpha_{3}\eta_{t}^{2})M_{\overline{E}}(x^{t}-x^{*})+\alpha_{4}\eta_{t}^{2}$
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\Bigr{)}$ (70)
Where $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq(8+2B)c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq A\rho_{2}L$. We now present the formal version of Theorem 3 as follows:
Theorem 9 (Formal version of Theorem 3).
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (B.3), if $x^{t}$ is generated by (51) with all assumptions in B.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\infty,\overline{E}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|x^{0}-x^{*}\|^{2}_{\infty,\overline{E}}+\frac{8\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{2c_{u}C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (71)
where $C_{1}=G(1+2x_{sp})$, $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $G$ is defined in (66) and $x_{sp}\coloneqq\sup\|x\|_{\infty,\overline{E}}$ is the upper bound of the span for all $x^{t}$.
Proof.
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (B.3) we have
$\mathbb{E}\bigl{[}M_{\overline{E}}(x^{t+1}-x^{*})|\mathcal{F}_{t}\Bigr{]}\leq(1-\alpha_{2}\eta_{t})M_{\overline{E}}(x^{t}-x^{*})+\alpha_{4}\eta_{t}^{2}+\eta_{t}C_{1}\varepsilon_{\text{bias}}$ (72)
we define $\Gamma_{t}\coloneqq\Pi_{i=o}^{t-1}(1-\alpha_{2}\eta_{t})$ and further obtain the $T$-step recursion relationship as follows:
$\mathbb{E}$ $\Bigl{[}M_{\overline{E}}(x^{T}-x^{*})\Bigr{]}\leq\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{4}\eta_{t}^{2}+\eta_{t}C_{1}\varepsilon_{\text{bias}}]$
$=\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{4}\eta_{t}^{2}]+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\eta_{t}C_{1}\varepsilon_{\text{bias}}]$
$=\underbrace{\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\frac{\alpha_{4}\Gamma_{T}}{\alpha_{2}}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{2}\eta_{t}^{2}]}_{R_{1}}+\underbrace{\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\eta_{t}C_{1}\varepsilon_{\text{bias}}]}_{R_{2}}$ (73)
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of (Zhang et al., 2021) which leads to
$R_{1}\leq\frac{K^{2}}{(T+K)^{2}}M_{\overline{E}}(x^{0}-x^{*})+\frac{4\alpha_{2}}{(T+K)\alpha_{2}^{2}}$ (74)
also, $R_{2}$ can be bounded by a logrithmic dependence of $T$
$R_{2}\leq\sum_{t=0}^{T-1}[\eta_{t}C_{1}\varepsilon_{\text{bias}}]=C_{1}\varepsilon_{\text{bias}}\sum_{t=0}^{T-1}\frac{1}{\alpha_{2}(t+K)}\leq\frac{C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (75)
Combining (74) and (75) with (73) would obtain the following:
$\mathbb{E}\Bigl{[}M_{\overline{E}}(x^{T}-x^{*})\Bigr{]}\leq\frac{K^{2}}{(T+K)^{2}}M_{\overline{E}}(x^{0}-x^{*})+\frac{4\alpha_{2}}{(T+K)\alpha_{2}^{2}}+\frac{C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (76)
Combining (76) with (52) yields (71). ∎
Appendix C Uncertainty Set Support Function Estimators
C.1 Proof of Theorem 4
We have
$\mathbb{E}[M]$ $=\sum_{n=0}^{N_{\max}-1}2^{n+1}\,\mathbb{P}(N^{\prime}=n)+2^{N_{\max}+1}\mathbb{P}(N^{\prime}=N_{\max})$
$=\sum_{n=0}^{N_{\max}-1}2^{n+1}\,\mathbb{P}(N=n)+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=\sum_{n=0}^{N_{\max}-1}\Bigl{(}\frac{2^{n+1}}{2^{n+1}}\Bigr{)}+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=N_{\max}+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=N_{\max}+\frac{2^{N_{\max}+1}}{2^{N_{\max}}}$
$=N_{\max}+2=\operatorname*{\mathcal{O}}(N_{\max}).$ (77)
C.2 Proof of Theorem 5
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untuncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From (Wang et al., 2023b), under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
$\mathbb{E}$ $\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]=\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]-\mathbb{E}\left[\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)\right]$
$=\mathbb{E}\left[\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right]-\mathbb{E}\left[\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N+1}}(V)+\frac{\Delta_{N}(V)}{\mathbb{P}(N=n)}\right]$
$=\mathbb{E}\left[\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right]-\mathbb{E}\left[\frac{\Delta_{N}(V)}{\mathbb{P}(N=n)}\right]$
$=\sum_{n=0}^{N_{\max}}\Delta_{n}(V)-\sum_{n=0}^{\infty}\Delta_{n}(V)$
$=\sum_{n=N_{\max}+1}^{\infty}\Delta_{n}(V)$ (78)
For each $\Delta_{n}(V)$, the expectation of absolute value can be bounded as
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]=\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]$
$+\frac{1}{2}\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a,E}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]+\frac{1}{2}\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a,O}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]$ (79)
By the binomial concentration and the Lipschitz property of the support function as in Lemma 2, we know for TV distance uncertainty, we have
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6(1+\frac{1}{\delta})2^{-\frac{n}{2}}\|V\|_{\mathrm{sp}}$ (80)
and for Wasserstein disance uncertainty, we have
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6\cdot 2^{-\frac{n}{2}}\|V\|_{\mathrm{sp}}$ (81)
Thus, for TV distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq\sum_{n=N_{\max}+1}^{\infty}\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (82)
and for Wasserstein distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq\sum_{n=N_{\max}+1}^{\infty}\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (83)
C.3 Proof of Lemma 2
For TV uncertainty sets, for a fixed $V$, for any $p\in\Delta(\mathcal{S})$, define $f_{p}(\mathbf{\mu})\coloneqq p(V-\mathbf{\mu})-\delta\|V-\mathbf{\mu}\|_{\mathrm{sp}}$ and $\mathbf{\mu}_{p}^{*}\coloneqq\arg\max_{\mathbf{\mu}\geq\mathbf{0}}f_{p}(\mathbf{\mu})$. Thus, we have
$\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)=f_{p}(\mu_{p}^{*})-f_{q}(\mu_{q}^{*})$ (84)
since, $\mathbf{\mu}_{p}^{*}$ and $\mathbf{\mu}_{q}^{*}$ are maximizers of $f_{p}$ and $f_{q}$ respectively, we further have
$f_{p}(\mathbf{\mu}_{q}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})\leq f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})\leq f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{p}^{*})$ (85)
Combing (84) and (85) we thus have:
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|$ $\leq\max\{|f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{p}^{*})|,|f_{p}(\mathbf{\mu}_{q}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})|\}$
$=\max\{|(p-q)(V-\mathbf{\mu}_{p}^{*})|,|(p-q)(V-\mathbf{\mu}_{q}^{*})|\}$ (86)
Note that $\sigma_{\mathcal{P}_{TV}}(V)$ can also be expressed as $\sigma_{\mathcal{P}_{TV}}(V)=p\mathbf{x}^{*}-\delta\|\mathbf{x}^{*}\|_{\mathrm{sp}}$ where $\mathbf{x}^{*}\coloneqq\arg\max_{\mathbf{x}\leq V}(p\mathbf{x}-\delta\|\mathbf{x}\|_{\mathrm{sp}})$. Let $M\coloneqq\max_{s}\mathbf{x}^{*}(s)$ and $m\coloneqq\min_{s}\mathbf{x}^{*}(s)$, then $\|\mathbf{x}\|_{\mathrm{sp}}=M-m$. Denote $\mathbf{e}$ as the all-ones vector, then $\mathbf{x}=\min_{s}V(s)\cdot\mathbf{e}$ is a feasible solution. Thus,
$p\mathbf{x}^{*}-\delta(M-m)\geq p(\min_{s}V(s)\cdot\mathbf{e})-\delta\|\min_{s}V(s)\cdot\mathbf{e}\|_{\mathrm{sp}}=\min_{s}V(s)$ (87)
Since $p$ is a probability vector, $p\mathbf{x}^{*}\leq M$, using the fact that $\delta>0$, we then obtain
$M-\delta(M-m)\geq\min_{s}V(s)\Rightarrow M-m\leq\frac{M-\min_{s}V(s)}{\delta}$ (88)
Since $\mathbf{x^{*}}$ is a feasible solution, we have
$M\leq\max_{s}V(s)\Rightarrow M-\min_{s}V(s)\leq\max_{s}V(s)-\min_{s}V(s)=\|V\|_{\mathrm{sp}}$ (89)
Combining (88) and (89) we obtain
$M-m\leq\frac{\|V\|_{\mathrm{sp}}}{\delta}\Rightarrow m\geq M-\frac{\|V\|_{\mathrm{sp}}}{\delta}\geq\min_{s}V(s)-\frac{\|V\|_{\mathrm{sp}}}{\delta}$ (90)
Where the last inequality is from $M\geq\min_{s}V(s)$, which is a direct result of (88) and the term $\delta(M-m)$ being positive. We finally arrive with
$\mathbf{x}^{*}(j)\in[m,M]\subseteq\Big{[}\min_{s}V(s)-\frac{\|V\|_{\mathrm{sp}}}{\delta},\max_{s}V(s)\Big{]}\quad\forall j\in\mathcal{S}$ (91)
Thus, $\|\mathbf{x}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}$, which leads to
$\|V-\mathbf{\mu}_{p}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}},\quad\|V-\mathbf{\mu}_{q}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}$ (92)
Combining (92) with (C.3) we obtain (28).
For Wasserstein uncertainty sets, note that for any $p\in\Delta(\mathcal{S})$ and value function $V$,
$\sigma_{\mathcal{P}_{W}}(V)=\sup_{\lambda\geq 0}\left(\overbrace{-\lambda\delta^{l}+\mathbb{E}_{p}\big{[}\underbrace{\inf_{y\in\mathcal{S}}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}}_{\phi(s,\lambda)}\big{]}}^{g(\lambda,p)}\right).$ (93)
Note that
$\inf_{y\in\mathcal{S}}V(y)\leq\phi(s,\lambda)\leq V(s)+\lambda d(S,s)^{l}=V(s)$ (94)
where the first inequality is because $\lambda d(S,y)^{l}\geq 0$ for any $d$ and $l$. We can then bound $\phi$ by the span of $V$ as
$|\phi(s,\lambda)|\leq\|V\|_{\mathrm{sp}}\quad\forall\lambda\geq 0$ (95)
We then further have that for any $p,q\in\Delta(\mathcal{S})$ and $\lambda\geq 0$,
$|g(\lambda,p)-g(\lambda,q)|\leq\sum_{s\in\mathcal{S}}|p(s)-q(s)||\phi(s,\lambda)|\leq\|p-q\|_{1}\|V\|_{\mathrm{sp}}$ (96)
using (96) and the fact that $|f(\lambda)-g(\lambda)|\leq\epsilon\Rightarrow|\sup_{\lambda}f(\lambda)-\sup_{\lambda}g(\lambda)|\leq\epsilon$, we obtain (29).
C.4 Proof of Theorem 6
For all $p\in\Delta(\mathcal{S})$, we have $\sigma_{p}(V)\leq\|V\|_{\mathrm{sp}}$, leading to
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]+\|V\|^{2}_{\mathrm{sp}}$ (97)
To bound the second moment, note that
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $=\mathbb{E}\left[\left(\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq\mathbb{E}\left[\left(\|V\|_{\mathrm{sp}}+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq 2\|V\|_{\mathrm{sp}}^{2}+2\mathbb{E}\left[\left(\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\left(\frac{\mathbb{E}[|\Delta_{n}(V)|]}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\mathbb{P}(N^{\prime}=n)$
$=2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{\mathbb{E}[|\Delta_{n}(V)|]^{2}}{\mathbb{P}(N^{\prime}=n)}$ (98)
Under TV distance uncertainty set, by (80), we further have
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{36(1+\frac{1}{\delta})^{2}2^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (99)
Under Wasserstein distance uncertainty set, by (81), we further have
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{362^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (100)
Appendix D Convergence for Robust TD
D.1 Formal Statement of Theorem 7
The first half of Algorithm 2 (line 1 - line 7) can be treated as a special instance of the SA updates in (51) with the bias and variance of the i.i.d. noise term specified in Section 6. We start with analyzing the bias and variance of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ for each $t$. Recall the definition of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is as follows:
$\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g_{0}+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\big{]}\quad\forall s\in\mathcal{S}$
Thus, we have for all $s\in\mathcal{S}$,
$\left|\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right|\leq\sum_{a}\pi(a|s)\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|=\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|$ (101)
Which further implies the bias of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is bounded by the bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ as follows:
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|$ (102)
Regarding the variance, note that
$\mathbb{E}\left[(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-{\mathbf{T}}_{g_{0}}(V_{t})(s))^{2}\right]$ $=\left(\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)^{2}+\mathrm{Var}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)$
$\leq\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|^{2}+\mathrm{Var}\left(\sum_{a}\pi(a|s)\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right)$
$=\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|^{2}+\sum_{a}\pi(a|s)^{2}\mathrm{Var}\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right)$ (103)
To create an upper bound of $\|V\|_{\mathrm{sp}}$ for all possible $V$, define the mixing time of any $p\in\mathcal{P}$ to be
$t^{p}_{\mathrm{mix}}\coloneqq\arg\min_{t\geq 1}\left\{\max_{\mu_{0}}\left\|(\mu_{0}p_{\pi}^{t})^{\top}-\nu^{\top}\right\|_{1}\leq\frac{1}{2}\right\}$ (104)
where $p_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. By Assumption 1, and Lemma 6,and for any value function $V$, we have
$t^{p}_{\mathrm{mix}}<+\infty\quad\text{and}\quad\|V\|_{\mathrm{sp}}\leq 4t^{p}_{\mathrm{mix}}$ (105)
Thus, define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma 5, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to (142) and combining the above result with Lemma 6, we obtian that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]\leq\|V_{t}\|^{2}\leq 16t^{2}_{\mathrm{mix}}$ (106)
Regarding TV distance uncertainty sets, using the property of the bias and variance of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ in Theorem 5 and Theorem 6 while combining them with Lemma 6, we have
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (107)
and
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$ (108)
Similarly, for Wasserstein distance uncertainty sets, we have
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (109)
and
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304t_{\mathrm{mix}}^{2}N_{\max}$ (110)
Thus, using the fact that the span is less or equal to two times the $l_{\infty}$ norm, line 1 - line 7 of Algorithm 2 can be treated as a special instance of the SA updates in (51) with $B=0$. Furthermore, for contamination uncertainty sets, we have
$\varepsilon^{\text{Cont}}_{\text{bias}}=0\quad\text{and}\quad A^{\text{Cont}}=16t^{2}_{\mathrm{mix}}$ (111)
for TV distance uncertainty sets, we have
$\varepsilon^{\text{TV}}_{\text{bias}}=48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (112)
$A^{\text{TV}}=2\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96t_{\mathrm{mix}}^{2}+4608(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (113)
and for Wasserstein distance uncertainty sets, we have
$\varepsilon^{\text{Wass}}_{\text{bias}}=48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (114)
$A^{\text{Wass}}=2\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96t_{\mathrm{mix}}^{2}+4608t_{\mathrm{mix}}^{2}N_{\max}=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (115)
Theorem 10 (Formal version of Theorem 7).
Let $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq 8c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq\rho_{2}L$, if $V_{t}$ is generated by Algorithm 2. Define $V^{*}$ to be the anchored robust value function $V^{*}=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$ for some $c$ such that $V^{*}(s_{0})=0$, then under Assumption 1 and if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$, then for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (116)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}$ (117)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (118)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}$ (119)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (120)
where $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8t_{\mathrm{mix}})$, $\gamma$ is defined in (17), $c_{u},c_{l}$ are defined in (52), $\rho_{2}$ is defined in (55) and $G$ is defined in (66).
Proof.
This is the result of substituting the terms of (111)-(115) and Theorem 2 to Theorem 9, and relating the span semi-norm to the $\|\cdot\|_{\infty,\overline{E}}$ semi-norm by (49). ∎
D.2 Proof of Theorem 7
We use the result from Theorem 10, to set $\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\epsilon^{2}$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, conbining with Theorem 4, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
D.3 Formal Statement of Theorem 8
To analyze the second part (line 8 - line 14) of Algorithm 2 and provide the provide the complexity for $g_{t}$, we first define the noiseless function $\bar{\delta}(V)$ as
$\bar{\delta}(V)\coloneqq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\big{[}r(s,a)+{\sigma}_{\mathcal{P}^{a}_{s}}(V)\big{]}-V(s)\right)$ (121)
Thus, we have
$\bar{\delta}_{t}=\bar{\delta}(V_{T})+\nu_{t}$ (122)
where $\nu_{t}$ is the noise term with bias equal to the bias $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})$
$\mathbb{E}[\left|\nu_{t}\right|]=\frac{1}{S}\sum_{s}\sum_{a}\left(\pi(a|s)\mathbb{E}\big{[}\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right|\big{]}\right)=\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|$ (123)
By the Bellman equation in Theorem 1, we have $g^{\pi}_{\mathcal{P}}=\bar{\delta}(V^{*})$, which implies
$\left|\bar{\delta}(V_{T})-g^{\pi}_{\mathcal{P}}\right|$ $=\left|\bar{\delta}(V_{T})-\bar{\delta}(V^{*})\right|$
$\leq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V^{*})\right|+\left|V_{T}(s)-V^{*}(s)\right|\right)$
$\leq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V^{*})\right|+\left|V_{T}(s)-V^{*}(s)\right|\right)$
$\leq\frac{1}{S}\sum_{s}2{\|V_{T}-V^{*}\|}_{\mathrm{sp}}$
$=2{\|V_{T}-V^{*}\|}_{\mathrm{sp}}$ (124)
Thus, the following recursion can be formed
$\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|$ $=\left|g_{t}+\beta_{t}(\bar{\delta}_{t}-g_{t})-g^{\pi}_{\mathcal{P}}\right|$
$=\left|g_{t}-g^{\pi}_{\mathcal{P}}+\beta_{t}(\bar{\delta}_{t}-g^{\pi}_{\mathcal{P}}+g^{\pi}_{\mathcal{P}}-g_{t})\right|$
$=\left|g_{t}-g^{\pi}_{\mathcal{P}}+\beta_{t}(\bar{\delta}(V^{T})-g^{\pi}_{\mathcal{P}}+\nu_{t}+g^{\pi}_{\mathcal{P}}-g_{t})\right|$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(\left|\bar{\delta}(V^{T})-g^{\pi}_{\mathcal{P}}\right|+\left|\nu_{t}\right|)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(2{\|V_{T}-V^{*}\|}_{\mathrm{sp}}+\left|\nu_{t}\right|)$ (125)
Thus, taking expectation conditioned on the filtration $\mathcal{F}^{t}$ yields
$\mathbb{E}\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]$ $\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (126)
By letting $\zeta_{t}\coloneqq\Pi_{i=0}^{t-1}(1-\beta_{t})$, we obtain the $T$-step recursion as follows:
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]$ $\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (127)
By setting $\beta_{t}\coloneqq\frac{1}{t+1}$, we have $\zeta_{T}=\frac{1}{T+1}\leq\frac{1}{T}$ and $\sum_{t=0}^{T-1}\beta_{t}\leq 2\log T$, (127) implies
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (128)
Theorem 11 (Formal version of Theorem 8).
Following all notations and assumptions in Theorem 10, then for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$ (129)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right)$ (130)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$
$+\frac{4\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$ (131)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (132)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$
$+\frac{4\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{\sqrt{\alpha_{2}}}+48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$ (133)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (134)
Proof.
By Theorem 10, taking square root on both side and utilizing the concavity of square root function, we have for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{Cont}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}$ (135)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{TV}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}}$ (136)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{Wass}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}}$ (137)
In addition, combining (82)-(83) with Lemma 6, we have for for TV distance uncertainty,
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (138)
and for Wasserstein distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (139)
Combining (135)-(139) with (128) gives the desired result. ∎
D.4 Proof of Theorem 8
We use the result from Theorem 11, to set $\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 4, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
Appendix E Some Auxiliary Lemmas for the Proofs
Lemma 4 (Lemma 6 in (Zhang et al., 2021)).
Under the setup and notation in Appendix B.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$, we have
$\mathbb{E}\left[\|x^{t+1}-x^{t}\|^{2}_{s,\overline{E}}|\mathcal{F}^{t}\right]\leq(16+4B)c_{u}\rho_{2}\eta_{t}^{2}M_{\overline{E}}(x^{t}-x^{*})+2A\rho_{2}\eta_{t}^{2}.$ (140)
Lemma 5 (Theorem D.1 in (Wang et al., 2023b)).
The estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (22) for contamination uncertainty sets is unbiased and has bounded variance as follows:
$\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]={\sigma}_{\mathcal{P}^{a}_{s}}(V),\quad\text{and}\quad\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\|V\|^{2}$ (141)
Lemma 6 (Ergodic case of Lemma 9 in (Wang et al., 2022)).
For any average reward MDP with stationary policy $\pi$ and the mixing time defined as
$\tau_{\mathrm{mix}}\coloneqq\arg\min_{t\geq 1}\left\{\max_{\mu_{0}}\left\|(\mu_{0}P_{\pi}^{t})^{\top}-\nu^{\top}\right\|_{1}\leq\frac{1}{2}\right\}$ (142)
where $P_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. If $P_{\pi}$ is irreducible and aperiodic, then $\tau_{\mathrm{mix}}<+\infty$ and for any value function $V$, we have
$\|V\|_{\mathrm{sp}}\leq 4\tau_{\mathrm{mix}}$ (143)