### v2

# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

Abstract
We present the first finite-sample analysis of policy evaluation in robust average-reward Markov Decision Processes (MDPs). Prior work in this setting have established only asymptotic convergence guarantees, leaving open the question of sample complexity. In this work, we address this gap by showing that the robust Bellman operator is a contraction under a carefully constructed semi-norm, and developing a stochastic approximation framework with controlled bias. Our approach builds upon Multi-Level Monte Carlo (MLMC) techniques to estimate the robust Bellman operator efficiently. To overcome the infinite expected sample complexity inherent in standard MLMC, we introduce a truncation mechanism based on a geometric distribution, ensuring a finite expected sample complexity while maintaining a small bias that decays exponentially with the truncation level. Our method achieves the order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for robust policy evaluation and robust average reward estimation, marking a significant advancement in robust reinforcement learning theory.
1 Introduction
Reinforcement learning (RL) has achieved notable success in domains such as robotics, finance, and healthcare, by enabling agents to learn optimal decision-making strategies through interaction with an environment. However, in many real-world applications, direct interaction is impractical because of safety concerns, high costs, or limited data‑collection budgets Sünderhauf et al. (2018); Höfer et al. (2021). This challenge is particularly evident in scenarios where agents are trained in simulated environments before being deployed in the real world, such as in robotic control and autonomous driving. The mismatch between simulated and real environments, known as the simulation-to-reality gap, often leads to performance degradation when the learned policy encounters unmodeled uncertainties. Robust reinforcement learning (robust RL) addresses this challenge by formulating the learning problem as an optimization over an uncertainty set of transition probabilities, ensuring reliable performance under worst-case conditions. In this work, we focus on the problem of evaluating the robust value function and robust average reward for a given policy using only data sampled from a simulator (nominal model), aiming to enhance generalization and mitigate the impact of transition uncertainty in real-world deployment.
Reinforcement learning problems under infinite time horizons are typically studied under two primary reward formulations: the discounted reward setting, where future rewards are exponentially discounted, and the average reward setting, which focuses on optimizing long-term performance. While the discounted-reward formulation is widely used, it may lead to myopic policies that underperform in applications requiring sustained long-term efficiency, such as queueing systems, inventory management, and network control. In contrast, the average-reward setting is more suitable for environments where decisions impact long-term operational efficiency. Despite its advantages, robust reinforcement learning under the average-reward criterion remains largely unexplored. Existing works on robust average-reward RL primarily provide asymptotic guarantees Wang et al. (2023a, b, 2024), lacking algorithms with finite-time performance bounds. This gap highlights the need for principled approaches that ensure robustness against model uncertainties while maintaining strong long-term performance guarantees.
Solving the robust average-reward reinforcement learning problem is significantly more challenging than its non-robust counterpart, with the primary difficulty arising in policy evaluation. Specifically, the goal is to compute the worst-case value function and worst-case average reward over an entire uncertainty set of transition models while having access only to samples from a nominal transition model. In this paper, we investigate three types of uncertainty sets: Contamination uncertainty sets, TV distance uncertainty sets, and Wasserstein distance uncertainty sets. Unlike the standard average-reward setting, where value functions and average rewards can be estimated directly from observed trajectories, the robust setting introduces an additional layer of complexity due to the need to optimize against adversarial transitions. Consequently, conventional approaches based on direct estimation such as Wei et al. (2020); Agarwal et al. (2021) immediately fail, as they do not account for the worst-case nature of the problem. Overcoming this challenge requires new algorithmic techniques that can infer the worst-case dynamics using only limited samples from the nominal model.
1.1 Challenges and Contributions
A common approach to policy evaluation in robust RL is to solve the corresponding robust Bellman operator. However, robust average-reward RL presents additional difficulties compared to the robust discounted setting. In the discounted case, the presence of a discount factor induces a contraction property in the robust Bellman operator Wang and Zou (2022); Zhou et al. (2024), facilitating stable iterative updates. In contrast, the average-reward Bellman operator lacks a contraction property with respect to any norm even in the non-robust setting Zhang et al. (2021), making standard fixed-point analysis inapplicable. Due to this fundamental limitation, existing works on robust average-reward RL such as Wang et al. (2023b) rely on asymptotic techniques, primarily leveraging ordinary differential equation (ODE) analysis to examine the behavior of temporal difference (TD) learning. These methods exploit the asymptotic stability of the corresponding ODE Borkar (2023) to establish almost sure convergence but fail to provide finite-sample performance guarantees. Addressing this limitation requires novel analytical tools and algorithmic techniques capable of providing explicit finite-sample bounds for robust policy evaluation and optimization.
In this work, we first establish and exploit a key structural property of the robust average-reward Bellman operator with uncertainty set $\mathcal{P}$ under the ergodicity assumption: its contraction under some semi-norm, denoted as $\|\cdot\|_{\mathcal{P}}$, where the detailed construction is specified in Theorem 2 and (15). Constructing $\|\cdot\|_{\mathcal{P}}$ is not straightforward, because ergodicity alone only guarantees that the chain mixes over multiple steps and fails to produce a single‐step contraction for familiar measures such as the span semi-norm. To overcome this, we group together all the worst‐case transition dynamics under uncertainty into one compact family of linear mappings, and observe that their “worst‐case gain” over any number of steps stays strictly below $1$. From this we build an extremal norm, which by construction shrinks every non‐constant component by the same fixed factor in a single step. Finally, we add a small “quotient” correction that exactly annihilates constant shifts, producing a semi-norm that vanishes only on constant functions but still inherits the one‐step shrinkage. The above construction yields a uniform, strict contraction for the robust Bellman operator.
This fundamental result above enables the use of stochastic approximation techniques similar to Zhang et al. (2021) to analyze and bound the error in policy evaluation, overcoming the lack of a standard contraction property that has hindered prior finite-sample analyses. Building on this insight, we develop a novel stochastic approximation framework tailored to the robust average-reward setting. Our approach simultaneously estimates both the robust value function and the robust average reward, leading to an efficient iterative procedure for solving the robust Bellman equation. A critical challenge in this framework under TV and Wasserstein distance uncertainty sets is accurately estimating the worst-case transition effects, which requires computing the support function of the uncertainty set. While previous works Blanchet and Glynn (2015); Blanchet et al. (2019); Wang et al. (2023b) have leveraged Multi-Level Monte Carlo (MLMC) for this task, their MLMC-based estimators suffer from infinite expected sample complexity due to the unbounded nature of the required geometric sampling, which lead to only asymptotic convergence. To address this, we introduce a truncation mechanism based on a truncated geometric distribution, ensuring that the sample complexity remains finite while maintaining an exponentially decaying bias. With these techniques, we derive the first finite-sample complexity guarantee for policy evaluation in robust average-reward RL, achieving an optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity bound. The main contributions of this paper are summarized as follows:
  • We prove that under the ergodicity assumption, the robust average-reward Bellman operator is a contraction with respect to a suitably constructed semi-norm (Theorem 2). This key result enables the application of stochastic approximation techniques for policy evaluation.
  • We prove the convergence of stochastic approximation under the semi-norm contraction and under i.i.d. with noise with non-zero bias (Theorem 8) as an intermediate result.
  • We develop an efficient method for computing estimates for the robust Bellman operator under TV distance and Wasserstein distance uncertainty sets. By modifying MLMC with a truncated geometric sampling scheme, we ensure finite expected sample complexity while keeping variance controlled and bias decaying exponentially with truncation level (Theorem 3-5).
  • We propose a novel temporal difference learning method that iteratively updates the robust value function and the robust average reward, facilitating efficient policy evaluation in robust average-reward RL. We establish the first non-asymptotic sample complexity result for policy evaluation in robust average-reward RL, proving an order-optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity for policy evaluation (Theorem 6), along with a $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity for robust average reward estimation (Theorem 7).
2 Formulation
2.1 Robust average-reward MDPs.
For a robust MDP with state space $\mathcal{S}$ and action space $\mathcal{A}$ while $|\mathcal{S}|=S$ and $|\mathcal{A}|=A$, the transition kernel is assumed to be in some uncertainty set $\mathcal{P}$. At each time step, the environment transits to the next state according to an arbitrary transition kernel $\mathsf{P}\in\mathcal{P}$. In this paper, we focus on the $(s,a)$-rectangular compact uncertainty set Nilim and Ghaoui (2003); Iyengar (2005), i.e., $\mathcal{P}=\bigotimes_{s,a}\mathcal{P}^{a}_{s}$, where $\mathcal{P}^{a}_{s}\subseteq\Delta(\mathcal{S})$. Popular uncertainty sets include those defined by the contamination model Huber (1965); Wang and Zou (2022), total variation Lim et al. (2013), and Wasserstein distance Gao and Kleywegt (2022).
We investigate the worst-case average-reward over the uncertainty set of MDPs. Specifically, define the robust average-reward of a policy $\pi$ as
$g^{\pi}_{\mathcal{P}}(s)\triangleq\min_{\kappa\in\bigotimes_{n\geq 0}\mathcal{P}}\lim_{T\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{T}\sum^{T-1}_{t=0}r_{t}|S_{0}=s\right],$ (1)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{n\geq 0}\mathcal{P}$. It was shown in Wang et al. (2023a) that the worst case under the time-varying model is equivalent to the one under the stationary model:
$g^{\pi}_{\mathcal{P}}(s)=\min_{\mathsf{P}\in\mathcal{P}}\lim_{T\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\left[\frac{1}{T}\sum^{T-1}_{t=0}r_{t}|S_{0}=s\right].$ (2)
Therefore, we limit our focus to the stationary model. We refer to the minimizers of (2) as the worst-case transition kernels for the policy $\pi$, and denote the set of all possible worst-case transition kernels by $\Omega^{\pi}_{g}$, i.e., $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$, where $g^{\pi}_{\mathsf{P}}$ denotes the average reward of policy $\pi$ under the single transition $\mathsf{P}\in\mathcal{P}$:
$g_{\mathsf{P}}^{\pi}(s)\triangleq\lim_{T\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}\frac{1}{T}\sum^{T-1}_{n=0}r_{t}|S_{0}=s\bigg{]}.$ (3)
We focus on the model-free setting, where only samples from the nominal MDP denoted as $\tilde{\mathsf{P}}$ (the centroid of the uncertainty set) are available. We investigate the problem of robust policy evaluation and robust average reward estimation, which means for a given policy $\pi$, we aim to estimate the robust value function and the robust average reward. We now formally define the robust value function $V^{\pi}_{\mathsf{P}_{V}}$ by connecting it with the following robust Bellman equation:
Theorem 1 (Robust Bellman Equation, Theorem 3.1 in Wang et al. (2023b)).
If $(g,V)$ is a solution to the robust Bellman equation
$V(s)=\sum_{a}\pi(a|s)\big{(}r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{)},\quad\forall s\in\mathcal{S},$ (4)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)=\min_{p\in\mathcal{P}^{a}_{s}}pV$ is denoted as the support function, then the scalar $g$ corresponds to the robust average reward, i.e., $g=g^{\pi}_{\mathcal{P}}$, and the worst-case transition kernel $\mathsf{P}_{V}$ belongs to the set of minimizing transition kernels, i.e., $\mathsf{P}_{V}\in\Omega^{\pi}_{g}$, where $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$. Furthermore, the function $V$ is unique up to an additive constant, where if $V$ is a solution to the Bellman equation, then we have $V=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$, where $c\in\mathbb{R}$ and $\mathbf{e}$ is the all-ones vector in $\mathbb{R}^{S}$, and $V^{\pi}_{\mathsf{P}_{V}}$ is definied as the relative value function of the policy $\pi$ under the single transition $\mathsf{P}_{V}$ as follows:
$V^{\pi}_{\mathsf{P}_{V}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}_{V}}\bigg{[}\sum^{\infty}_{t=0}(r_{t}-g^{\pi}_{\mathsf{P}_{V}})|S_{0}=s\bigg{]}.$ (5)
Theorem 1 implies that the robust Bellman equation (4) identifies both the worst‐case average reward $g$ and a corresponding value function $V$ that is determined only up to an additive constant. In particular, $\sigma_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$. Unlike the robust discounted case, where the contraction property of the Bellman operator under the sup-norm enables straightforward fixed-point iteration, the robust average-reward Bellman equation does not induce contraction under any norm, making direct iterative methods inapplicable. Throughout the paper, we denote $\mathbf{e}$ as the all-ones vector in $\mathbb{R}^{S}$. We now characterize the explicit forms of $\sigma_{\mathcal{P}^{a}_{s}}(V)$ for different compact uncertainty sets are as follows:
Contamination Uncertainty Set The $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Under this uncertainty set, the support function can be computed as
$\sigma_{\mathcal{P}^{a}_{s}}(V)=(1-\delta)\tilde{\mathsf{P}}^{a}_{s}V+\delta\min_{s}V(s),$ (6)
and this is linear in the nominal transition kernel $\mathsf{P}^{a}_{s}$.
Total Variation Uncertainty Set. The total variation uncertainty set is $\mathcal{P}^{a}_{s}=\{q\in\Delta(|\mathcal{S}|):\frac{1}{2}\|q-\tilde{\mathsf{P}}^{a}_{s}\|_{1}\leq\delta\},$ define $\|\cdot\|_{\mathrm{sp}}$ as the span semi-norm and the support function can be computed using its dual function Iyengar (2005):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\max_{\mu\geq\mathbf{0}}\big{(}\tilde{\mathsf{P}}^{a}_{s}(V-\mu)-\delta\|V-\mu\|_{\mathrm{sp}}\big{)}.$ (7)
Wasserstein Distance Uncertainty Sets. Consider the metric space $(\mathcal{S},d)$ by defining some distance metric $d$. For some parameter $l\in[1,\infty)$ and two distributions $p,q\in\Delta(\mathcal{S})$, define the $l$-Wasserstein distance between them as $W_{l}(q,p)=\inf_{\mu\in\Gamma(p,q)}\|d\|_{\mu,l}$, where $\Gamma(p,q)$ denotes the distributions over $\mathcal{S}\times\mathcal{S}$ with marginal distributions $p,q$, and $\|d\|_{\mu,l}=\big{(}\mathbb{E}_{(X,Y)\sim\mu}\big{[}d(X,Y)^{l}\big{]}\big{)}^{1/l}$. The Wasserstein distance uncertainty set is then defined as
$\mathcal{P}^{a}_{s}=\left\{q\in\Delta(|\mathcal{S}|):W_{l}(\tilde{\mathsf{P}}^{a}_{s},q)\leq\delta\right\}.$ (8)
The support function w.r.t. the Wasserstein distance set, can be calculated as follows Gao and Kleywegt (2023):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\sup_{\lambda\geq 0}\left(-\lambda\delta^{l}+\mathbb{E}_{\tilde{\mathsf{P}}^{a}_{s}}\big{[}\inf_{y}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}\big{]}\right).$ (9)
2.2 Robust Bellman Operator
Motivated by Theorem 1, we define the robust Bellman operator, which forms the basis for our policy evaluation procedure.
Definition 1 (Robust Bellman Operator, Wang et al. (2023b)).
The robust Bellman operator $\mathbf{T}_{g}$ is defined as:
$\mathbf{T}_{g}(V)(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{]},\quad\forall s\in\mathcal{S}.$ (10)
The operator $\mathbf{T}_{g}$ transforms a value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under some constructed semi-norm. This allows us to develop provably efficient stochastic approximation algorithms. Throughout this paper, we make the following standard assumption regarding the structure of the induced Markov chain.
Assumption 1.
The Markov chain induced by $\pi$ is irreducible and aperiodic for all $\mathsf{P}\in\mathcal{P}$.
Assumption 1 is used widely in the current robust average reinforcement learning literature Wang et al. (2023a, b, 2024); Sun et al. (2024). This assumption ensures that, under any transition model within the uncertainty set, the policy $\pi$ induces a single recurrent communicating class. A well-known result in average-reward MDPs states that under Assumption 1, the average reward is independent of the starting state, i.e., for any $\mathsf{P}\in\mathcal{P}$ and all $s,s^{\prime}\in\mathcal{S}$, we have $g^{\pi}_{\mathsf{P}}(s)=g^{\pi}_{\mathsf{P}}(s^{\prime})$. Thus, we can drop the dependence on the initial state and simply write $g^{\pi}_{\mathsf{P}}$ as the robust average reward.
Remark 1.
We note that when the radius of uncertainty sets are small enough, Assumption 1 could be relaxed into assuming only the nominal transition being irreducible and aperiodic. We provide the detailed discussions in Appendix B.
3 Semi-Norm Contraction of Robust Bellman Operators
Under Assumption 1, we are able to establish the semi-norm contraction property. For motivation, we first establish the semi-norm contraction property of the non-robust average reward Bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows:
$\mathbf{T}_{g}^{\mathsf{P}}(V)(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g+\sum_{s^{\prime}}\mathsf{P}(s^{\prime}|s,a)V(s^{\prime})\big{]},\quad\forall s\in\mathcal{S}.$ (11)
Lemma 1.
Let $\mathcal{S}$ be a finite state space, and let $\pi$ be a stationary policy. If the Markov chain induced by $\pi$ under the transition $\mathsf{P}$ is irreducible and aperiodic, there exists a semi-norm $\|\cdot\|_{\mathsf{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ and a constant $\beta\in(0,1)$ such that for all $V_{1},V_{2}\in\mathbb{R}^{S}$ and any $g\in\mathbb{R}$,
$\bigl{\|}\mathbf{T}_{g}^{\mathsf{P}}(V_{1})-\mathbf{T}_{g}^{\mathsf{P}}(V_{2})\bigr{\|}_{\mathsf{P}}\leq\beta\|V_{1}-V_{2}\|_{\mathsf{P}}.$ (12)
Proof Sketch
Under ergodicity, the one‐step transition matrix (denoted as $\mathsf{P}^{\pi}$) has a unique stationary distribution $d^{\pi}$, define the stationary projector $E=\mathbf{e}^{\top}d^{\pi}$, then the fluctuation matrix (defined as $Q^{\pi}=\mathsf{P}^{\pi}-E$) has all eigenvalues strictly inside the unit circle. Standard finite‐dimensional theory (via the discrete Lyapunov equation Horn and Johnson (2012)) would produce a norm $\|\cdot\|_{Q}$ on $\mathbb{R}^{S}$ such that there is a constant $\alpha\in(0,1)$ such that for any $x\in\mathbb{R}^{S}$, $\|Q^{\pi}x\|_{Q}\leq\alpha\|x\|_{Q}$. We then build the semi-norm as follows:
$\|x\|_{\mathsf{P}}=\|Q^{\pi}x\|_{Q}+\epsilon\inf_{c\in\mathbb{R}}\|x-c\mathbf{e}\|_{Q},\quad 0<\epsilon<1-\alpha,$ (13)
so that its kernel is exactly the constant vectors (the second term vanishes only on shifts of $\mathbf{e}$) and the first term enforces a one‐step shrinkage by $\beta=\alpha+\epsilon<1$. A short calculation then shows $\|\mathsf{P}^{\pi}x\|_{\mathsf{P}}\leq\beta\|x\|_{\mathsf{P}}$, yielding the desired contraction, which leads to the overall result.
The concrete proof of Lemma 1 including the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$ is in Appendix C.1, where the properties of irreducible and aperiodic finite state Markov chain are utilized. Thus, we show the (non-robust) average reward Bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under $\|\cdot\|_{\mathsf{P}}$. Based on the above motivations, we now formally establish the contraction property of the robust average reward Bellman operator by leveraging Lemma 1 and the compactness of the uncertainty sets.
Theorem 2.
Under Assumption 1, and if $\mathcal{P}$ is compact, there exists a semi-norm $\|\cdot\|_{\mathcal{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ such that the robust Bellman operator $\mathbf{T}_{g}$ is a contraction. Specifically, there exist $\gamma\in(0,1)$ such that
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathcal{P}}\;\leq\;\gamma\,\|V_{1}-V_{2}\|_{\mathcal{P}},\;\forall\;V_{1},V_{2}\in\mathbb{R}^{S},\,g\in\mathbb{R}.$ (14)
Proof Sketch
For any $\mathsf{P}\in\mathcal{P}$, the one‐step transition matrix $\mathsf{P}^{\pi}$ has a unique stationary projector $E_{\mathsf{P}}$ due to ergodicity. Since $\mathcal{P}$ is compact, the family of fluctuation matrices $\{Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ has joint spectral radius strictly less than $1$. By Lemma 10in Berger and Wang (1992), one is able to construct an “extremal norm” (denoted as $\|\cdot\|_{\rm ext}$) under which every $Q_{\mathsf{P}}^{\pi}$ contracts by a uniform factor $\alpha\in(0,1)$. Mimicking the non-robust case in Lemma 1, we similarly define
$\|x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\|x-c\mathbf{e}\|_{\rm ext},\quad 0<\epsilon<1-\alpha.$ (15)
The supremum term zeros out if $x\in\{c\mathbf{e}:c\in\mathbb{R}\}$ , and it inherits the uniform one‐step shrinkage by $\alpha$. Adding the small quotient term fixes the kernel without spoiling $\gamma=\alpha+\epsilon<1$, so one shows at once
$\|\mathbf{T}^{\mathsf{P}}_{g}(V_{1})-\mathbf{T}^{\mathsf{P}}_{g}(V_{2})\|_{\mathcal{P}}\leq\gamma\|V_{1}-V_{2}\|_{\mathcal{P}}\quad\text{for all}\;\mathsf{P}\in\mathcal{P}$ (16)
The above leads to the desired results.
The concrete proof of Theorem 2 along with the detailed construction of the semi-norm $\|\cdot\|_{\mathcal{P}}$ are in Appendix C.2. Since all the uncertainty sets listed in Section 2.1 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the contraction property in Theorem 2.
4 Efficient Estimators for Uncertainty Sets
To utilize the contraction property in Section 3 to obtain convergence rate results, our idea is perform the following iterative stochastic approximation:
$V_{t+1}(s)\leftarrow V_{t}(s)+\eta_{t}\left(\hat{\mathbf{T}}_{g}(V_{t})(s)-V_{t}(s)\right),\quad\forall s\in\mathcal{S}$ (17)
where the learning rate $\eta_{t}$ would be specified in Section 5. The detailed analysis and complexities of the general stochastic approximation in the form of (17) is provided in Appendix D. Theorem 8 implies that if $\hat{\mathbf{T}}_{g}(V)$, being an estimator of $\mathbf{T}_{g}(V)$, could be constructed with bounded variance and small bias, $V_{t}$ converges to a solution of the Bellman equation in (4). However, the challenge of constructing our desired $\hat{\mathbf{T}}_{g}(V)$ lies in the construction of the support function estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$.
In this section, we aim to construct an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ in various uncertainty sets. Recall that the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$ as defined in the robust Bellman equation in Theorem 1. The explicit forms of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for different uncertainty sets were characterized in (6)-(9). Our goal in this section is to construct efficient estimators $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ that approximates ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ while maintaining controlled variance and finite sample complexity.
Linear Contamination Uncertainty Set
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Since the support function can be computed by (6) and the expression is linear in the nominal transition kernel $\tilde{\mathsf{P}}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
$\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\triangleq(1-\delta)V(s^{\prime})+\delta\min_{x}V(x),$ (18)
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (18) is just one. Lemma 12 from Wang et al. (2023b) states that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (18) is unbiased and has bounded variance as follows:
$\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]={\sigma}_{\mathcal{P}^{a}_{s}}(V),\quad\text{and}\quad\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\|V\|^{2}$ (19)
Nonlinear Contamination Sets
Regarding TV and Wasserstein distance uncertainty sets, they have a nonlinear relationship between the nominal distribution $\tilde{\mathsf{P}}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (Blanchet and Glynn, 2015; Blanchet et al., 2019; Wang et al., 2023b) have proposed a multi-level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approaches require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator as $\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$. To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a truncated-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
Algorithm 1 Truncated MLMC Estimator for TV and Wasserstein Uncertainty Sets
Input: $s\in\mathcal{S}$, $a\in\mathcal{A}$, Max level $N_{\max}$, Value function $V$
1:Sample $N\sim\mathrm{Geom}(0.5)$
2:$N^{\prime}\leftarrow\min\{N,N_{\max}\}$
3:Collect $2^{N^{\prime}+1}$ i.i.d. samples of $\{s^{\prime}_{i}\}^{2^{N^{\prime}+1}}_{i=1}$ with $s^{\prime}_{i}\sim\tilde{\mathsf{P}}^{a}_{s}$ for each $i$
4:$\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i}\}}$
5:$\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}}}\sum_{i=1}^{2^{N^{\prime}}}\mathbbm{1}_{\{s^{\prime}_{2i-1}\}}$
6:$\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}\leftarrow\frac{1}{2^{N^{\prime}+1}}\sum_{i=1}^{2^{N^{\prime}+1}}\mathbbm{1}_{\{s^{\prime}_{i}\}}$
7:$\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}\leftarrow\mathbbm{1}_{\{s^{\prime}_{1}\}}$
8:if TV then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (7)
9:else if Wasserstein then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (9)
10:$\Delta_{N^{\prime}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V)-\frac{1}{2}\Bigl{[}\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V)+\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)\Bigr{]}$
11:$\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\leftarrow\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)},\text{where }p^{\prime}(n)=\mathbb{P}(N^{\prime}=n)$ return $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$
In particular, if $n<N_{\max}$, then $\{N^{\prime}=n\}=\{N=n\}$ with probability $(\tfrac{1}{2})^{n+1}$, while $\{N^{\prime}=N_{\max}\}$ has probability $\sum_{m=N_{\max}}^{\infty}(1/2)^{m+1}=2^{-N_{\max}}$. After obtaining $N^{\prime}$, Algorithm 1 then collects a set of $2^{N^{\prime}+1}$ i.i.d. samples from the nominal transition model to construct empirical estimators for different transition distributions. The core of the approach lies in computing the support function estimates for TV and Wasserstein uncertainty sets using a correction term $\Delta_{N^{\prime}}(V)$, which accounts for the bias introduced by truncation. This correction ensures that the final estimator maintains a low bias while achieving a finite sample complexity. We now present several crucial properties of Algorithm 1.
Theorem 3 (Finite Sample Complexity).
Under Algorithm 1, denote $M=2^{N^{\prime}+1}$ as the random number of samples (where $N^{\prime}=\min\{N,N_{\max}\}$). Then
$\mathbb{E}[M]=N_{\max}+2=\mathcal{O}(N_{\max}).$ (20)
The proof of Theorem 3 is in Appendix E.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
Theorem 4 (Exponentially Decaying Bias).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 the under TV uncertainty set, we have:
$\left|\mathbb{E}\bigl{[}\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\bigr{]}\right|\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (21)
where $\delta$ denotes the radius of TV distance. Under Wasserstein uncertainty set, we have:
$\left|\mathbb{E}\bigl{[}\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\bigr{]}\right|\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (22)
Theorem 4 establishes that the bias of the truncated MLMC estimator decays exponentially with $N_{\max}$, ensuring that truncation does not significantly affect accuracy. This result follows from observing that the deviation introduced by truncation can be expressed as a sum of differences between support function estimates at different level, and each of which is controlled by the $\ell_{1}$-distance between transition distributions. Thus, we can use binomial concentration property to ensure the exponentially decaying bias.
The proof of Theorem 4 is in Appendix E.2. One important lemma used in the proof is the following Lemma 2, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
Lemma 2.
For any $p,q\in\Delta(\mathcal{S})$, let $\mathcal{P}_{TV}$ and $\mathcal{Q}_{TV}$ denote the TV distance uncertainty set with radius $\delta$ centering at $p$ and $q$ respectively, and let $\mathcal{P}_{W}$ and $\mathcal{Q}_{W}$ denote the Wasserstein distance uncertainty set with radius $\delta$ centering at $p$ and $q$ respectively. Then for any value function $V$, we have:
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}\;\text{and}\;|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (23)
We refer the proof of Theorem 4 to Appendix E.3.
Theorem 5 (Linear Variance).
Let $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ be the estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained from Algorithm 1 then under TV distance uncertainty set, we have:
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq 3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (24)
and under Wasserstein distance uncertainty set, we have:
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq 3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (25)
Theorem 5 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 5 is in Appendix E.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
5 Robust Average Reward TD Learning
Equipped with the methods of constructing $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we now present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section D, ensuring both efficient sample complexity and finite-time convergence guarantees.
The algorithm is divided into two main phases. The first phase (Lines 1-7) estimates the robust value function. The noisy Bellman operator is computed using the estimator $\hat{\sigma}_{\mathcal{P}_{s}^{a}}(V_{t})$ obtained depending on the uncertainty set type. Then the iterative update follows a stochastic approximation scheme with step size $\eta_{t}$, ensuring convergence while maintaining stability. Finally, the value function is centered at an anchor state $s_{0}$ to remove the ambiguity due to its additive invariance. The second phase (Lines 8-14) estimates the robust average reward by utilizing $V_{T}$ from the output of the first phase. The expected Bellman residual $\delta_{t}(s)$ is computed across all states and averaging it to obtain $\bar{\delta}_{t}$. A separate stochastic approximation update with step size $\beta_{t}$ is then applied to refine $g_{t}$, ensuring convergence to the robust worst-case average reward. By combining these two phases, Algorithm 2 provides an efficient and provably convergent method for robust policy evaluation under average-reward criteria, marking a significant advancement over prior methods that only provided asymptotic guarantees.
Algorithm 2 Robust Average Reward TD
Input: Policy $\pi$, Initial values $V_{0}$, $g_{0}=0$, Stepsizes $\eta_{t}$, $\beta_{t}$, Max level $N_{\max}$, Anchor state $s_{0}\in\mathcal{S}$
1:for $t=0,1,\ldots,T-1$ do
2:     for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
3:         if Contamination then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (18)
4:         else if TV or Wasserstein then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
5:     $\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\leftarrow\sum_{a}\pi(a|s)\big{[}r(s,a)-g_{0}+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\big{]},\quad\forall s\in\mathcal{S}$
6:     $V_{t+1}(s)\leftarrow V_{t}(s)+\eta_{t}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-V_{t}(s)\right),\quad\forall s\in\mathcal{S}$
7:     $V_{t+1}(s)=V_{t+1}(s)-V_{t+1}(s_{0}),\quad\forall s\in\mathcal{S}$
8:for $t=0,1,\ldots,T-1$ do
9:     for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
10:         if Contamination then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to (18)
11:         else if TV or Wasserstein then Sample $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ according to Algorithm 1
12:     $\hat{\delta}_{t}(s)\leftarrow\sum_{a}\pi(a|s)\big{[}r(s,a)+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\big{]}-V_{T}(s),\quad\forall s\in\mathcal{S}$
13:     $\bar{\delta}_{t}\leftarrow\frac{1}{S}\sum_{s}\hat{\delta}_{t}(s)$
14:     $g_{t+1}\leftarrow g_{t}+\beta_{t}(\bar{\delta}_{t}-g_{t})$return $V_{T}$, $g_{T}$
To derive the sample complexity of robust policy evaluation, we utilize the semi-norm contraction property of the Bellman operator in Theorem 2, and fit Algorithm 2 into the general biased stochastic approximation result in Theorem 8 while incorporating the bias analysis characterized in Section 4. Since each phase of Algorithm 2 contains a loop of length $T$ with all the states and actions updated together, the total samples needed for the entire algorithm in expectation is $2SAT\mathbb{E}[N_{\max}]$, where $\mathbb{E}[N_{\max}]$ is one for contamination uncertainty sets and is $\operatorname*{\mathcal{O}}(N_{\max})$ from Theorem 3 for TV and Wasserstein distance uncertainty sets.
Theorem 6.
If $V_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for contamination uncertainty set and a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for TV and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $V_{T}$. Moreover, these results are order-optimal in terms of $\epsilon$.
Theorem 7.
If $g_{t}$ is generated by Algorithm 2 and satisfying Assumption 1, then if the stepsize $\beta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$, we require a sample complexity of $\tilde{\operatorname*{\mathcal{O}}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ for all contamination, TV, and Wasserstein distance uncertainty set to ensure an $\epsilon$ convergence of $g_{T}$.
The formal version of Theorems 6 and 7 along with the proofs are in Appendix F. Theorem 6 provides the order-optimal sample complexity of $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ for Algorithm 2 to achieve an $\epsilon$-accurate estimate of $V_{T}$. The proof of Theorem 7 extends the analysis of Theorem 6 to robust average reward estimation. The key difficulty lies in controlling the propagation of error from value function estimates to reward estimation. By again leveraging the contraction property and appropriately tuning step sizes, we establish an $\tilde{\operatorname*{\mathcal{O}}}(\epsilon^{-2})$ complexity bound for robust average reward estimation.
6 Conclusion
This paper provides the first finite-sample analysis for policy evaluation in robust average-reward MDPs, bridging a gap where only asymptotic guarantees existed. By introducing a biased stochastic approximation framework and leveraging the properties of various uncertainty sets, we establish finite-time convergence under biased noise. Our algorithm achieves an order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for policy evaluation, despite the added complexity of robustness.
A crucial step in our analysis is proving that the robust Bellman operator is contractive under our constructed semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring the validity of stochastic approximation updates. We further develop a truncated multi-level Monte Carlo estimator that efficiently computes worst-case value functions under total variation and Wasserstein uncertainty, while keeping bias and variance controlled. One limitation of this work is that the results require ergodicity to hold in the setting, as stated in Assumption 1. Additionally, scaling the algorithm and results in the paper via function approximations remains an important open problem.
7 Acknowledgement
We would like to thank Zaiwei Chen for assisting in identifying relevant literature and providing constructive feedback.
Appendix A Related Work
The theoretical guarantees of robust average reward have been studied by the following works. Wang et al. (2023a) takes a model-based perspective, approximating robust average-reward MDPs with discounted MDPs and proving uniform convergence of the robust discounted value function as the discount factor approaches one, employing dynamic programming and Blackwell optimality arguments to characterize optimal policies. Wang et al. (2023b) propose a model-free approach by developing robust relative value iteration (RVI) TD and Q-learning algorithms, proving their almost sure convergence using stochastic approximation, martingale theory, and multi-level Monte Carlo estimators to handle non-linearity in the robust Bellman operator. While these studies provide fundamental insights into robust average-reward RL, they do not establish explicit convergence rate guarantees due to the lack of contraction properties in the robust Bellman operator. In addition, Sun et al. (2024) studies the policy optimization of average reward robust MDPs with assuming direct queries of the value functions.
Policy evaluation in robust discounted-reward reinforcement learning with finite sample guarantees has been extensively studied, with the key recent works Wang and Zou (2022); Zhou et al. (2024); Li et al. (2022); Kumar et al. (2023); Kuang et al. (2022) focusing on solving the robust Bellman equation by finding its fixed-point solution. This approach is made feasible by the contraction property of the robust Bellman operator under the sup-norm, which arises due to the presence of a discount factor $\gamma<1$. However, this fundamental approach does not directly extend to the robust average-reward setting, where the absence of a discount factor removes the contraction property under any norm. As a result, existing robust discounted methods cannot be applied in the robust average reward RL setting.
Appendix B Relaxing Assumption 1
In this section, we provide further justifications on Remark 1 and characterize the conditions when Assumption 1 could be relaxed into only assuming the nominal transition being irreducible and aperiodic for contamination, TV distance and Wasserstein distance uncertainty sets.
Lemma 3.
Let $\pi$ be a stationary policy on a finite state space $S$, and let $P_{0}$ be the nominal kernel under $\pi$ whose induced chain is irreducible and aperiodic. Define
$m_{\min}=\min_{\begin{subarray}{c}s,a,s^{\prime}\\ P_{0}(s^{\prime}|s,a)>0\end{subarray}}P_{0}(s^{\prime}|s,a)>0,\qquad d_{\min}=\min_{x\neq y}d(x,y)>0$
where $d$ is the metric for the Wasserstein uncertainty set. Then there exists $\bar{\delta}>0$ such that for every $\delta\in[0,\bar{\delta})$, any $P$ in each of the following families remains irreducible and aperiodic under $\pi$:
$\mathcal{P}_{\rm cont}(\delta)=\Bigl{\{}P:P=(1-\delta)P_{0}+\delta\,Q,\;Q\in\Delta(S)\Bigr{\}},$
$\mathcal{P}_{\rm TV}(\delta)=\Bigl{\{}Q\in\Delta(S):\tfrac{1}{2}\|Q-P_{0}\|_{1}\leq\delta\Bigr{\}},$
$\mathcal{P}_{W}(\delta)=\Bigl{\{}Q\in\Delta(S):W_{1}(Q,P_{0})\leq\delta\Bigr{\}}.$
Moreover one may take
$\bar{\delta}=\min\Bigl{\{}1,\;\tfrac{1}{2}\,m_{\min},\;m_{\min}\,d_{\min}\Bigr{\}}.$
Proof.
We show that for each uncertainty type, whenever $\delta<\bar{\delta}$ every originally positive transition under $P_{0}$ remains strictly positive in $P$. Preservation of the support–graph and any self–loops then guarantees irreducibility and period one.
1. Contamination. If
$P(s^{\prime}|s,a)=(1-\delta)\,P_{0}(s^{\prime}|s,a)+\delta\,Q(s^{\prime}|s,a),$
then
$P(s^{\prime}\!\mid s,a)\geq(1-\delta)P_{0}(s^{\prime}\!\mid s,a).$
Since $P_{0}(s^{\prime}\!\mid s,a)>0$ and $\delta<1$, each positive edge of $P_{0}$ survives, as do any self–loops.
2. Total‐variation. If $\tfrac{1}{2}\|Q-P_{0}\|_{1}\leq\delta$, then
$|Q(s^{\prime})-P_{0}(s^{\prime})|\leq 2\delta,\quad\Longrightarrow\quad Q(s^{\prime})\geq P_{0}(s^{\prime})-2\delta.$
Define $m_{\min}=\min\{P_{0}(s^{\prime}\!\mid s,a)>0\}$. For $\delta<\tfrac{1}{2}\,m_{\min}$, we have $Q(s^{\prime}\!\mid s,a)\geq m_{\min}-2\delta>0$. Thus all original transitions remain.
3. Wasserstein. If $W_{1}(Q,P_{0})\leq\delta$, moving mass $P_{0}(s^{\prime})$ from $s^{\prime}$ to any other state costs at least
$P_{0}(s^{\prime})\,d_{\min}.$
Hence $Q(s^{\prime})=0$ would force $W_{1}(Q,P_{0})\geq P_{0}(s^{\prime})\,d_{\min}$. Taking $\delta<m_{\min}\,d_{\min}$ prevents zeroing any originally positive entry.
In all three cases, the perturbed kernel $P$ retains exactly the positive‐support and self–loops of $P_{0}$. Therefore irreducibility and aperiodicity hold uniformly over $\mathcal{P}_{\rm cont}(\delta)$, $\mathcal{P}_{\rm TV}(\delta)$, and $\mathcal{P}_{W}(\delta)$ whenever $\delta<\bar{\delta}$. ∎
Appendix C Semi-Norm Contraction Property of the Bellman Operator
C.1 Proof of Lemma 1
For any $V_{1},V_{2}\in\mathbb{R}^{S}$ and define $\Delta=V_{1}-V_{2}$. Denote $\mathsf{P}^{\pi}$ as the transition matrix under policy $\pi$ and the unique stationary distribution $d^{\pi}$, and denote $E$ as the matrix with all rows being identical to $d^{\pi}$. We further define $Q^{\pi}=\mathsf{P}^{\pi}-E$. Thus, we would have,
$\mathbf{T}_{g}^{\pi}(V_{1})(s)-\mathbf{T}_{g}^{\pi}(V_{2})(s)=\sum_{s^{\prime}\in\mathcal{S}}\mathsf{P}^{\pi}(s^{\prime}|s)\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}=\mathsf{P}^{\pi}\,\Delta(s).$ (26)
which implies
$\mathbf{T}_{g}^{\pi}(V_{1})-\mathbf{T}_{g}^{\pi}(V_{2})=\mathsf{P}^{\pi}\Delta=Q^{\pi}\Delta+E\Delta$ (27)
We now discuss the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$. Since $\mathsf{P}^{\pi}$ is ergodic, according to the Perron–Frobenius theorem, $\mathsf{P}^{\pi}$ has an eigenvalue $\lambda_{1}=1$ of algebraic multiplicity exactly one, with corresponding right eigenvector $\mathbf{e}$. Moreover, all other eigenvalues $\lambda_{2}\geq\ldots\geq\lambda_{S}$ of $\mathsf{P}^{\pi}$ satisfies $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$.
Lemma 4.
All eigenvalues of $Q^{\pi}$ lies strictly inside the unit circle.
Proof.
Since $E=\mathbf{e}d^{\pi\top}$, $E$ is a rank‑one projector onto the span of $\mathbf{e}$. Hence the spectrum of $E$ is $\{1,0,\ldots,0\}$. In addition, we can show $\mathsf{P}^{\pi}$ and $E$ commute by
$\mathsf{P}^{\pi}E=\mathsf{P}^{\pi}(\mathbf{e}d^{\pi})=(\mathbf{e}d^{\pi})=E$ (28)
$E\mathsf{P}^{\pi}=\mathbf{e}(d^{\pi}\mathsf{P}^{\pi})=\mathbf{e}d^{\pi}=E$ (29)
Thus, by the Schur theorem, $\mathsf{P}$ and $E$ are simultaneously upper triangularizable. In a common triangular basis, the diagonals of $\mathsf{P}$ and $E$ list their eigenvalues in descending orders, which are $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{S}\}$ and $\{1,0,\ldots,0\}$ respectively. Thus, in that same basis, $Q^{\pi}=\mathsf{P}^{\pi}-E$ is also triangular, with diagonal entries being $\{\lambda_{1}-1,\lambda_{2}-0,\ldots,\lambda_{S}-0\}$. Since $\lambda_{1}=1$, we have the spectrum of $Q^{\pi}$ is exactly $\{\lambda_{2},\ldots,\lambda_{S},0\}$. Since we already have $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$, we conclude the proof. ∎
Define $\rho(\cdot)$ to be the spectral radius of a matrix, then Lemma 4 implies that $\rho(Q^{\pi})<1$. Hence by equivalence of norms in $\mathbb{R}^{|\mathcal{S}|}$ it is possible to construct a vector norm $\|\cdot\|_{Q}$ so that the induced operator norm of $Q^{\pi}$ is less than $1$, specifically
$10\leq\rho(Q^{\pi})\leq\bigl{\|}Q^{\pi}\bigr{\|}_{Q}\leq\alpha<$ 1 . (30)
A concrete construction example is to leverage the discrete-Lyapunov equation Horn and Johnson (2012) of solving $M$ on the space of symmetric matrices for any $\rho(Q^{\pi})<\alpha<1$ as follows:
$Q^{\pi\top}MQ^{\pi}-\alpha^{2}M=-I$ (31)
We can express $M$ in the form of Neumann‐series as
$M$ $=\alpha^{-2}I+\alpha^{-4}(Q^{\pi})^{\top}Q^{\pi}+\alpha^{-6}\left((Q^{\pi})^{\top}\right)^{2}(Q^{\pi})^{2}+\dots$
$=\sum_{k=0}^{\infty}\alpha^{-2(k+1)}\left((Q^{\pi})^{\top}\right)^{k}(Q^{\pi})^{k}.$ (32)
Since $\rho((Q^{\pi})^{\top}Q^{\pi})=\rho(Q^{\pi})^{2}<\alpha^{2}<1$, the partial sum expression in (C.1) converges absolutely as a geometric‐type series. Also, since each term in (C.1) is positive semi‑definite, and the first term $\alpha^{-2}I$ being positive definite, we can conclude that $M$ being the summation is well-defined and is positive definite. Thus, using the $M$ defined in (31), we can define our desired $\|\cdot\|_{Q}$ as
$\|x\|_{Q}\coloneqq\sqrt{x^{\top}Mx}$ (33)
which implies
$\|Q^{\pi}\|_{Q}=\sup_{x\neq\bf{0}}\frac{\|Q^{\pi}x\|_{Q}}{\|x\|_{Q}}=\sup_{x\neq\bf{0}}\frac{\sqrt{(Q^{\pi}x)^{\top}MQ^{\pi}x}}{\sqrt{x^{\top}Mx}}\overset{(a)}{\leq}\alpha<1$ (34)
Where $(a)$ is because for any $x\neq\bf{0}$, from (31) we have
$(Qx)^{\pi\top}MQ^{\pi}x-\alpha^{2}x^{\top}Mx=-x^{\top}x\quad\Rightarrow\quad(Qx)^{\pi\top}MQ^{\pi}x=\alpha^{2}x^{\top}Mx-\|x\|_{2}^{2}$ (35)
Since $\|x\|_{2}^{2}$ is always non-negative dividing both sides of the second equation of (35) by $x^{\top}Mx$ and further taking the square root on both sides yields the inequality of $(a)$.
Based on the above construction of $\|\cdot\|_{Q}$, define the operator $\|\cdot\|_{\mathsf{P}}$ as
$\|x\|_{\mathsf{P}}\coloneqq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{Q}$ (36)
where $0<\epsilon<1-\alpha$.
Lemma 5.
The operator $\|\cdot\|_{\mathsf{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathsf{P}}\leq(\alpha+\epsilon)\bigl{\|}x\bigr{\|}_{\mathsf{P}}$.
Proof.
Regarding positive homogeneity and nonnegativity, for any scalar $\lambda$ and $x\in\mathbb{R}^{S}$,
$\|\lambda x\|_{\mathsf{P}}=\bigl{\|}Q^{\pi}(\lambda x)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}\lambda x-c\mathbf{e}\bigr{\|}_{Q}=|\lambda|\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\epsilon\,|\lambda|\inf_{c}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{Q}=|\lambda|\,\|x\|_{\mathsf{P}},$
and clearly $\|x\|_{\mathsf{P}}\geq 0$, with equality only when both $\|Q^{\pi}x\|_{Q}=0$ and $\inf_{c}\|x-c\,e\|_{Q}=0$. Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$,
$\|x+y\|_{\mathsf{P}}$ $=\bigl{\|}Q^{\pi}(x+y)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}x+y-c\mathbf{e}\bigr{\|}_{Q}$
$\leq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\bigl{\|}Q^{\pi}y\bigr{\|}_{Q}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{Q}$
$\leq\bigl{\|}Q^{\pi}x\bigr{\|}_{Q}+\bigl{\|}Q^{\pi}y\bigr{\|}_{Q}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{Q}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{Q}$
$=\|x\|_{\mathsf{P}}+\|y\|_{\mathsf{P}}.$
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then we have
$\|x\|_{\mathsf{P}}$ $=\|kQ^{\pi}\mathbf{e}\|_{Q}+\epsilon\inf_{c}\|k\mathbf{e}-c\mathbf{e}\|_{Q}$
$=\|k(\mathsf{P}^{\pi}-E)\mathbf{e}\|_{Q}+\epsilon\|k\mathbf{e}-k\mathbf{e}\|_{Q}$
$=\|k\mathbf{e}-k\mathbf{e}\|_{Q}+\epsilon\|0\|_{Q}=0$ (37)
On the other hand, if $x\notin\{c\mathbf{e}:c\in\mathbb{R}\}$, we know that
$\|x\|_{\mathsf{P}}\geq\epsilon\inf_{c}\|x-c\mathbf{e}\|_{Q}>0$ (38)
Thus, the kernel of $\|\cdot\|_{\mathsf{P}}$ is exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. We now show that, for any $x\in\mathbb{R}^{S}$,
$\bigl{\|}Q^{\pi}x\bigr{\|}_{\mathsf{P}}$ $=\bigl{\|}Q^{\pi}(Q^{\pi}x)\bigr{\|}_{Q}+\epsilon\inf_{c}\bigl{\|}Q^{\pi}x-c\mathbf{e}\bigr{\|}_{Q}$
$\leq\alpha\|Q^{\pi}x\|_{Q}+\epsilon\|Q^{\pi}x\|_{Q}$
$=(\alpha+\epsilon)\|Q^{\pi}x\|_{Q}$
$\leq(\alpha+\epsilon)\|x\|_{\mathsf{P}}.$ (39)
Let $\beta=\alpha+\epsilon$, by (30) and (36), we have $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, $\beta\in(0,1)$ and combining $\beta$ with the semi-norm $\|\cdot\|_{\mathsf{P}}$ confirms Lemma 1.
C.2 Proof of Theorem 2
We override the terms $\alpha,\lambda$ and $\epsilon$ from the previous section. For any $V_{1},V_{2}$ and $s\in\mathcal{S}$,
$\mathbf{T}_{g}(V_{1})(s)-$ $\mathbf{T}_{g}(V_{2})(s)=\sum_{a\in\mathcal{A}}\pi(a|s)[\sigma_{p^{a}_{s}}(V_{1})-\sigma_{p^{a}_{s}}(V_{2})]$
$=\sum_{a\in\mathcal{A}}\pi(a|s)[\min_{p\in\mathcal{P}^{a}_{s}}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\min_{p\in\mathcal{P}^{a}_{s}}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\max_{p\in\mathcal{P}^{a}_{s}}\Bigg{[}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})\Bigg{]}$
$\leq\sum_{a}\pi(a|s)\sum_{s^{\prime}}\tilde{p}_{(V_{1},V_{2})}(s^{\prime}|s,a)\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}$ (40)
where $\tilde{p}_{(V_{1},V_{2})}=\arg\max_{p\in\mathcal{P}^{a}_{s}}[\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{1}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)V_{2}(s^{\prime})]$ and each $\tilde{p}_{(V_{1},V_{2})}\in\mathcal{P}$ for all $V_{1},V_{2}$. We now discuss the construction of the desired semi-norm $\|\cdot\|_{\mathcal{P}}$.
For any $\mathsf{P}\in\mathcal{P}$, denote $\mathsf{P}^{\pi}$ as the transition matrix under policy $\pi$ and the unique stationary distribution $d^{\pi}_{\mathsf{P}}$, and denote $E_{\mathsf{P}}$ as the matrix with all rows being identical to $d^{\pi}_{\mathsf{P}}$. We further define $Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}$. Thus, by Lemma 4 we have $\rho(Q_{\mathsf{P}}^{\pi})<1$ for all $\mathsf{P}\in\mathcal{P}$. Since $\mathcal{P}$ is compact, denote $r^{*}$ as the supremum of the spectral radius of all $Q_{\mathsf{P}}^{\pi}$, we have:
$r^{*}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\rho(Q_{\mathsf{P}}^{\pi})<1$ (41)
We follow similar process for constructing our desired semi-norm $\|\cdot\|_{\mathcal{P}}$ as in Appendix C.1 by first constructing a norm such that all $Q_{\mathsf{P}}^{\pi}$ are strictly less then one under that norm. We choose $\alpha\in(r^{*},1)$ and we follow the approach in Wirth (2002) by constructing an extremal norm $\|\cdot\|_{\rm ext}$ as follows:
$\|x\|_{\rm ext}\coloneqq\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\quad\text{where}\quad Q^{\pi}_{\mathcal{P}}=\{Q^{\pi}_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ (42)
Note that we follow the convention that $\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}=\|x\|_{2}$ when $k=0$.
Lemma 6.
The operator $\|\cdot\|_{\rm ext}$ is a valid norm with $\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}<1$ for all $\mathsf{P}\in\mathcal{P}$
Proof.
We first prove that $\|\cdot\|_{\rm ext}$ is bounded. Following Lemma 10 and choosing $\lambda\in(r^{*},\alpha)$, then there exist a positive constant $C<\infty$ such that
$\|Q_{k}Q_{k-1}\ldots Q_{1}\|_{2}\leq C\lambda^{k}$ (43)
Hence for each $k$ and for all $x\in\mathbb{R}^{S}$,
$\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\leq\alpha^{-k}C\lambda^{k}\|x\|_{2}=C\left(\frac{\lambda}{\alpha}\right)^{-k}\|x\|_{0}\longrightarrow 0\quad\text{as}\quad k\rightarrow\infty$ (44)
Thus the double supremum in (42) is over a bounded and vanishing sequence, so $\|\cdot\|_{\rm ext}$ bounded.
To check that $\|\cdot\|_{\rm ext}$ is a valid norm, note that if $x=\bf{0}$, $\|x\|_{\rm ext}$ is directly 00. On the other hand, if $\|x\|_{\rm ext}=0$, we have
$\sup_{k=0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}=\|x\|_{2}=0\Rightarrow x=\bf{0}$ (45)
Regarding homogeneity, observe that for any $c\in\mathbb{R}$ and $x\in\mathbb{R}^{S}$,
$\|cx\|_{\rm ext}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(cx)\|_{2}=|c|\|x\|_{\rm ext}$ (46)
Regarding triangle inequality, using $\|Q_{k}\ldots Q_{1}(x+y)\|_{2}\leq\|Q_{k}\ldots Q_{1}x\|_{2}+\|Q_{k}\ldots Q_{1}y\|_{2}$ for any $x,y\in\mathbb{R}^{S}$, we obtain,
$\|x+y\|_{\rm est}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(x+y)\|_{2}\leq\|x\|_{\rm est}+\|y\|_{\rm est}$ (47)
For any $\mathsf{P}\in\mathcal{P}$, we have
$\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}$ $=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(Q_{\mathsf{P}}^{\pi}x)\|_{2}$
$\leq\sup_{k\geq 1}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-(k-1)}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$=\alpha\sup_{k\geq 1}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$\leq\alpha\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}$
$=\alpha\|x\|_{\rm ext}$ (48)
Since $\mathsf{P}$ is arbitrary, (6) implies that for any $\mathsf{P}\in\mathcal{P}$,
$\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}=\sup_{x\neq\bf{0}}\frac{\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}}{\|x\|_{\rm ext}}\leq\alpha<1$ (49)
We now follow the same method as (36) to construct the semi-norm $\|\cdot\|_{\mathcal{P}}$. Define the operator $\|\cdot\|_{\mathcal{P}}$ as
$\|x\|_{\mathcal{P}}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}$ (50)
where $0<\epsilon<1-\alpha$.
Lemma 7.
The operator $\|\cdot\|_{\mathcal{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathcal{P}}\leq(\alpha+\epsilon)\bigl{\|}x\bigr{\|}_{\mathcal{P}}$.
Proof.
Regarding positive homogeneity and nonnegativity, for any scalar $\lambda$ and $x\in\mathbb{R}^{S}$,
$\|\lambda x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(\lambda x)\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}\lambda x-c\mathbf{e}\bigr{\|}_{\rm ext}=|\lambda|\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon|\lambda|\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}=|\lambda|\|x\|_{\rm ext}$
and $\|x\|_{\rm ext}\geq 0$. Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$, note that for any $\mathsf{P}\in\mathcal{P}$,
$\|Q^{\pi}_{\mathsf{P}}(x+y)\|_{\rm ext}\leq\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}$ (51)
Taking supremum over $\mathsf{P}$ on both sides yields
$\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}(x+y)\|_{\rm ext}\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}$ (52)
Thus, we have
$\|x+y\|_{\mathcal{P}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(x+y)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}x+y-c\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{\rm ext}$
$=\|x\|_{\mathcal{P}}+\|y\|_{\mathcal{P}}.$
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then similar to (5), we have
$\|x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(k\mathbf{e})\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\|k\mathbf{e}-c\mathbf{e}\|_{\rm ext}=0+0=0$ (53)
On the other hand, if $x\notin\{c\mathbf{e}:c\in\mathbb{R}\}$, we know that
$\|x\|_{\mathcal{P}}\geq\epsilon\inf_{c}\|x-c\mathbf{e}\|_{\rm ext}>0$ (54)
Thus, the kernel of $\|\cdot\|_{\mathcal{P}}$ is exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. We now show that, for any $x\in\mathbb{R}^{S}$ and $\mathsf{P}\in\mathcal{P}$,
$\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathcal{P}}$ $=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}Q\mathsf{P}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}\mathsf{P}^{\pi}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}QQ_{\mathsf{P}}^{\pi}x+QE_{\mathsf{P}}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}Q_{\mathsf{P}}^{\pi}x+E_{\mathsf{P}}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$=\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}QQ_{\mathsf{P}}^{\pi}x\|_{\rm ext}+\epsilon\inf_{c}\bigl{\|}Q_{\mathsf{P}}^{\pi}x-c\mathbf{e}\bigr{\|}_{\rm ext}$
$\leq\alpha\sup_{Q\in\ Q_{\mathcal{P}}^{\pi}}\bigl{\|}Qx\|_{\rm ext}+\epsilon\|Q^{\pi}x\|_{\rm ext}$
$=(\alpha+\epsilon)\|Q^{\pi}x\|_{\rm ext}$
$\leq(\alpha+\epsilon)\|x\|_{\mathcal{P}}.$ (55)
Since $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, let $\gamma=\alpha+\epsilon$ then $\gamma\in(0,1)$. Substituting the above result back to (C.2), we obtain
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathcal{P}}\leq\|\pi(a|s)\tilde{p}_{(V_{1},V_{2})}\bigl{[}V_{1}(s^{\prime})-V_{2}(s^{\prime})\bigr{]}\|_{\mathcal{P}}\leq\gamma\|V_{1}-V_{2}\|_{\mathcal{P}}$ (56)
Appendix D Biased Stochastic Approximation Convergence Rate
In Section 3, we established that the robust Bellman operator is a contraction under the semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 4). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which is further used to establish the sample complexity of policy evaluation in robust average reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
$x^{t+1}=x^{t}+\eta_{t}\bigl{[}\widehat{H}(x^{t})-x^{t}\bigr{]},\quad\text{where}\quad\widehat{H}(x^{t})=H(x^{t})+w^{t}.$ (57)
with $\eta_{t}>0$ being the step-size sequence. We assume that there exist $\gamma\in(0,1)$ such that
$\|H(x)-H(y)\|_{\mathcal{P}}\leq\gamma\,\|x-y\|_{\mathcal{P}},\quad\forall x,y$ (58)
We also assume that the noise terms $\omega^{t}$ are i.i.d. and have bounded bias and variance
$\mathbb{E}[\,\|w^{t}\|_{\mathcal{P}}^{2}|\mathcal{F}^{t}]\leq A+B\,\|x^{t}-x^{*}\|_{\mathcal{P}}^{2}\quad\text{and}\quad\bigl{\|}\mathbb{E}[\,w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{P}}\leq\varepsilon_{\mathrm{bias}}$ (59)
Theorem 8.
If $x^{t}$ is generated by (57) with all assumptions in (58) and (59) satisfied, then if the stepsize $\eta_{t}\coloneqq\operatorname*{\mathcal{O}}(\frac{1}{t})$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathcal{P}}\Bigr{]}\leq\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}\right)\|x^{0}-x^{*}\|^{2}_{\mathcal{P}}+\operatorname*{\mathcal{O}}\left(\frac{A}{(1-\gamma)^{2}T}\right)+\operatorname*{\mathcal{O}}\left(\frac{x_{\mathrm{sup}}\varepsilon_{\text{bias}}\log T}{1-\gamma}\right)$ (60)
where $x_{\mathrm{sup}}\coloneqq\sup_{x}\|x\|_{\mathcal{P}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ semi-norm for all $x^{t}$.
Theorem 8 adapts the analysis of (Zhang et al., 2021) and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$.
We perform analysis of the biased‐noise extension to the semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift Zhang et al. (2021); Chen et al. (2025) and using properties in dual norms for managing the bias.
D.1 Proof of Theorem 8
D.1.1 Setup and Notation.
In this section, we override the notation of the semi-norm $\|\cdot\|_{\mathcal{P}}$ by re-writing it as the norm $\|\cdot\|_{\mathcal{N}}$ (defined in (64)) to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalent class $\delta_{\overline{E}}$, define the indicator function $\delta_{\overline{E}}$ as
$\delta_{\bar{E}}(x):=\begin{cases}0&x\in\bar{E},\\ \infty&\text{{otherwise}.}\end{cases}$ (61)
then by Zhang et al. (2021), the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalent class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
$\|x\|_{c,\overline{E}}\coloneqq(\|\cdot\|_{c}\ast_{\inf}\delta_{\overline{E}})(x)=\inf_{y}(\|x-y\|_{c}+\delta_{\overline{E}}(y))=\inf_{e\in\overline{E}}\|x-e\|_{c}\quad\forall x.$ (62)
Where $\ast_{\inf}$ denotes the infimal convolution operator. Throughout the remaining section, we let $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. Since $\|\cdot\|_{\mathcal{P}}$ constructed in (50) is a semi-norm with kernel being $\overline{E}$, we can construct a norm $\|\cdot\|_{\mathcal{N}}$ such that
$\|x\|_{\mathcal{N},\overline{E}}\coloneqq(\|\cdot\|_{\mathcal{N}}\ast_{\inf}\delta_{\overline{E}})(x)=\|x\|_{\mathcal{P}}$ (63)
We construct $\|\cdot\|_{\mathcal{N}}$ as follows:
$\|x\|_{\mathcal{N}}\coloneqq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}$ (64)
where $Q_{\mathsf{P}}^{\pi}$ and $\epsilon$ are defined in (50).
Lemma 8.
The operator $\|\cdot\|_{\mathcal{N}}$ defined in (64) is a norm satisfying (63).
Proof.
We first verify that $\|\cdot\|_{\mathcal{N}}$ is a norm. Regarding positivity, since all terms in (64) are non-negative, $\|x\|_{\mathcal{N}}\geq 0$ for all $x\in\mathbb{R}^{S}$ and $\|\mathbf{0}\|_{\mathcal{N}}=0$. If $x\neq\bf{0}$, since $\|\cdot\|_{\rm ext}$ is a valid norm and $\epsilon>0$, we have
$\|x\|_{\mathcal{N}}\geq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}>0.$
Regarding homogeneity, For any $\lambda\in\mathbb{R}$, we have
$\|\lambda x\|_{\mathcal{N}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(\lambda x)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}(\lambda x)-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|\lambda x\|_{\rm ext}$
$=|\lambda|\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon|\lambda|\inf_{c}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon|\lambda|\|x\|_{\rm ext}$
$=|\lambda|\|x\|_{\mathcal{N}}$
Regarding triangle inequality, for any $x,y\in\mathbb{R}^{S}$, we have
$\|x+y\|_{\mathcal{N}}$ $=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}(x+y)\bigr{\|}_{\rm ext}+\epsilon\inf_{c}\bigl{\|}(x+y)-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x+y\|_{\rm ext}$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a,b}\bigl{\|}x-a\mathbf{e}+y-b\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon(\|x\|_{\rm ext}+\|y\|_{\rm ext})$
$\leq\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}x\|_{\rm ext}+\sup_{\mathsf{P}\in\mathcal{P}}\|Q^{\pi}_{\mathsf{P}}y\|_{\rm ext}+\epsilon\inf_{a}\bigl{\|}x-a\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{b}\bigl{\|}y-b\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}+\|y\|_{\rm ext}$
$=\|x\|_{\mathcal{N}}+\|y\|_{\mathcal{N}}.$
We now show that since $Q^{\pi}_{\mathsf{P}}\mathbf{e}=0$ for all $\mathsf{P}\in\mathcal{P}$, by the definition of infimal convolution, we have that for all $x\in\mathbb{R}^{S}$,
$(\|\cdot\|_{\mathcal{N}}\ast_{\inf}\delta_{\overline{E}})(x)$ $=\inf_{k\in\mathbb{R}}\|x-k\mathbf{e}\|_{\mathcal{N}}$
$=\inf_{k\in\mathbb{R}}\left(\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x-kQ_{\mathsf{P}}^{\pi}\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}-k\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\inf_{k\in\mathbb{R}}\left(\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\bigl{\|}x-c\mathbf{e}\bigr{\|}_{\rm ext}+\inf_{k\in\mathbb{R}}\left(\epsilon\|x-k\mathbf{e}\|_{\rm ext}\right)$
$=\|x\|_{\mathcal{P}}$
We thus restate our problem of analyzing the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$, with the operator $H:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying the contraction property as follows:
$\|H(x)-H(y)\|_{\mathcal{N},\overline{E}}\leq\gamma\|x-y\|_{\mathcal{N},\overline{E}},\quad\gamma\in(0,1),\quad\forall x,y$ (65)
The stochastic approximation iteration being used is as follows
$x^{t+1}=x^{t}+\eta_{t}\bigl{[}\widehat{H}(x^{t})-x^{t}\bigr{]},\quad\text{where}\quad\widehat{H}(x^{t})=H(x^{t})+w^{t}.$ (66)
We assume:
  • $\mathbb{E}[\,\|w^{t}\|_{\mathcal{N},\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}^{2}$ (In the robust average reward TD case, $B=0$).
  • $\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{N},\overline{E}}\leq\varepsilon_{\mathrm{bias}}$.
  • $\eta_{t}>0$ is a chosen step‐size sequence (decreasing or constant).
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of Zhang et al. (2021). Thus, we emphasize mostly on managing the bias.
D.1.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By (Zhang et al., 2021, Proposition 1–2), using the Moreau envelope function $M(x)$ in Definition 2.2 of Chen et al. (2020), we define
$M_{\overline{E}}(x)=\bigl{(}M\ast_{\inf}\delta_{\overline{E}}\bigr{)}(x),$
so that there exist $c_{l},c_{u}>0$ with
$c_{l}M_{\overline{E}}(x)\leq\frac{1}{2}\|x\|_{\mathcal{N},\overline{E}}^{2}\leq c_{u}M_{\overline{E}}(x),$ (67)
and $M_{\overline{E}}$ is $L$-smooth w.r.t. another semi-norm $\|\cdot\|_{s,\overline{E}}$. Concretely, $L$-smoothness means:
$M_{\overline{E}}(y)\leq M_{\overline{E}}(x)+\langle\nabla M_{\overline{E}}(x),\,y-x\rangle+\tfrac{L}{2}\,\|\,y-x\|_{s,\overline{E}}^{2},\quad\forall\,x,y.$ (68)
Moreover, the gradient of $M_{\overline{E}}$ satisfies $\langle\nabla M_{\overline{E}}(x),\;c\,\mathbf{e}\rangle=0$ for all $x$, and the dual norm denoted as $\|\cdot\|_{*,s,\overline{E}}$ is also L-smooth:
$\|\nabla M_{\overline{E}}(x)-\nabla M_{\overline{E}}(y)\|_{*,s,\overline{E}}\leq L\|\,y-x\|_{s,\overline{E}},\quad\forall\,x,y.$ (69)
Note that since $\|\cdot\|_{s,\overline{E}}$ and $\|\cdot\|_{\mathcal{N},\overline{E}}$ are semi-norms on a finite‐dimensional space with the same kernel, there exist $\rho_{1},\rho_{2}>0$ such that
$\rho_{1}\,\|z\|_{\mathcal{N},\overline{E}}\leq\|z\|_{s,\overline{E}}\leq\rho_{2}\,\|z\|_{\mathcal{N},\overline{E}},\;\forall\,z.$ (70)
Likewise, their dual norms (denoted $\|\cdot\|_{*,s,\overline{E}}$ and $\|\cdot\|_{*,\mathcal{N},\overline{E}}$) satisfy the following:
$\frac{1}{\rho_{2}}\|z\|_{*,s,\overline{E}}\leq\|z\|_{*,\mathcal{N},\overline{E}}\leq\frac{1}{\rho_{1}}\|z\|_{*,s,\overline{E}},\;\forall\,z.$ (71)
D.1.3 Formal Statement of Theorem 8
By $L$‐smoothness w.r.t. $\|\cdot\|_{s,\overline{E}}$ in (68), for each $t$,
$M_{\overline{E}}(x^{t+1}-x^{*})\leq M_{\overline{E}}(x^{t}-x^{*})+\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\bigr{\rangle}+\tfrac{L}{2}\|\,x^{t+1}-x^{t}\|_{s,\overline{E}}^{2}.$ (72)
where $x^{t+1}-x^{t}=\eta_{t}[\widehat{H}(x^{t})-x^{t}]=\eta_{t}[H(x^{t})+w^{t}-x^{t}]$. Taking expectation of the second term of the RHS of (72) conditioned on the filtration $\mathcal{F}^{t}$ we obtain,
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),$ $x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]=\eta_{t}\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}+\omega^{t}\rangle|\mathcal{F}^{t}]$
$=\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\omega^{t}\rangle|\mathcal{F}^{t}]$
$=\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[\omega^{t}|\mathcal{F}^{t}]\rangle.$ (73)
[…]
$\bigl{\langle}u,\,v\bigr{\rangle}\leq\|u\|_{*}\,\|v\|,\quad\forall\,u,v.$ (74)
In the biased noise setting, $u=\nabla M_{\overline{E}}(x^{t}-x^{*})$ and $v=\mathbb{E}[w^{t}|\mathcal{F}^{t}]$, with $\|\cdot\|=\|\cdot\|_{\mathcal{N},\overline{E}}$. So
$\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq\bigl{\|}\nabla M_{\overline{E}}(x^{t}-x^{*})\bigr{\|}_{*,\,\mathcal{N},\overline{E}}\cdot\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{N},\overline{E}}.$ (75)
Since $\|\mathbb{E}[w^{t}|\mathcal{F}^{t}]\|_{\mathcal{N},\overline{E}}\leq\varepsilon_{\text{bias}}$, it remains to bound $\|\nabla M_{\overline{E}}(x^{t}-x^{*})\|_{*,\mathcal{N},\overline{E}}$. By setting $y=0$ in (69), we get
$\|\nabla M_{\overline{E}}(x)-\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}\leq L\|x\|_{s,\overline{E}},\quad\forall\,x.$ (76)
Thus,
$\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}\leq\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}},\quad\forall\,x.$ (77)
By (71), we know that there exists $\frac{1}{\rho_{2}}\leq\alpha\leq\frac{1}{\rho_{1}}$ such that
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}$ (78)
Thus, combining (77) and (78) would give:
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}}\big{)},\quad\forall\,x.$ (79)
By (70), we know that $\|x\|_{s,\overline{E}}\leq\|x\|_{\mathcal{N},\overline{E}}$, thus we have:
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\rho_{2}\|x\|_{\mathcal{N},\overline{E}}\big{)},\quad\forall\,x.$ (80)
Hence, combining the above with (75), there exist some
$G=\mathcal{O}\big{(}\frac{1}{\rho_{1}}\max\{L\rho_{2},\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}\}\big{)}$ (81)
such that
$\mathbb{E}\Bigl{[}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\,w^{t}\rangle|\mathcal{F}^{t}\Bigr{]}=\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq G\bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\bigr{)}\varepsilon_{\text{bias}}.$ (82)
Combining (82) with (D.1.3) we obtain
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]\leq\eta_{t}$ $\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle$
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\Bigr{)}$ (83)
To bound the first term in the RHS of (D.1.3), note that
$\langle\nabla M_{\bar{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle$ $=\langle\nabla M_{\bar{E}}(x^{t}-x^{*}),H(x^{t})-x^{*}+x^{*}-x^{t}\rangle$
$\overset{(a)}{\leq}M_{\bar{E}}(H(x^{t})-x^{*})-M_{\bar{E}}(x^{t}-x^{*})$
$\overset{(b)}{\leq}\frac{1}{2c_{l}}\|H(x^{t})-H(x^{*})\|^{2}_{c,\bar{E}}-M_{\bar{E}}(x^{t}-x^{*})$
$\overset{(c)}{\leq}\frac{\gamma^{2}}{2c_{l}}\|x^{t}-x^{*}\|^{2}_{c,\bar{E}}-M_{\bar{E}}(x^{t}-x^{*})$
$\leq\left(\frac{\gamma^{2}c_{u}}{c_{l}}-1\right)M_{\bar{E}}(x^{t}-x^{*})$
$\leq-(1-\gamma\sqrt{c_{u}/c_{l}})M_{\bar{E}}(x^{t}-x^{*}),$ (84)
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalent class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (D.1.3). (D.1.3) and Lemma 11 with (72), we arrive as follows:
$\mathbb{E}\Bigl{[}M_{\overline{E}}(x^{t+1}-x^{*})|\mathcal{F}_{t}\Bigr{]}$ $\leq(1-2\alpha_{2}\eta_{t}+\alpha_{3}\eta_{t}^{2})M_{\overline{E}}(x^{t}-x^{*})+\alpha_{4}\eta_{t}^{2}$
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\Bigr{)}$ (85)
Where $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq(8+2B)c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq A\rho_{2}L$. We now present the formal version of Theorem 8 as follows:
Theorem 9 (Formal version of Theorem 8).
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (D.1.3), if $x^{t}$ is generated by (66) with all assumptions in D.1.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathcal{N},\overline{E}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|x^{0}-x^{*}\|^{2}_{\mathcal{N},\overline{E}}+\frac{8\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{2c_{u}C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (86)
where $C_{1}=G(1+2x_{sup})$, $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $G$ is defined in (81) and $x_{sup}\coloneqq\sup\|x\|_{\mathcal{N},\overline{E}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ norm for all $x^{t}$.
Proof.
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (D.1.3) we have
$\mathbb{E}\bigl{[}M_{\overline{E}}(x^{t+1}-x^{*})|\mathcal{F}_{t}\Bigr{]}\leq(1-\alpha_{2}\eta_{t})M_{\overline{E}}(x^{t}-x^{*})+\alpha_{4}\eta_{t}^{2}+\eta_{t}C_{1}\varepsilon_{\text{bias}}$ (87)
we define $\Gamma_{t}\coloneqq\Pi_{i=o}^{t-1}(1-\alpha_{2}\eta_{t})$ and further obtain the $T$-step recursion relationship as follows:
$\mathbb{E}$ $\Bigl{[}M_{\overline{E}}(x^{T}-x^{*})\Bigr{]}\leq\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{4}\eta_{t}^{2}+\eta_{t}C_{1}\varepsilon_{\text{bias}}]$
$=\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{4}\eta_{t}^{2}]+\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\eta_{t}C_{1}\varepsilon_{\text{bias}}]$
$=\underbrace{\Gamma_{T}M_{\overline{E}}(x^{0}-x^{*})+\frac{\alpha_{4}\Gamma_{T}}{\alpha_{2}}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\alpha_{2}\eta_{t}^{2}]}_{R_{1}}+\underbrace{\Gamma_{T}\sum_{t=0}^{T-1}(\frac{1}{\Gamma_{t+1}})[\eta_{t}C_{1}\varepsilon_{\text{bias}}]}_{R_{2}}$ (88)
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of Zhang et al. (2021) which leads to
$R_{1}\leq\frac{K^{2}}{(T+K)^{2}}M_{\overline{E}}(x^{0}-x^{*})+\frac{4\alpha_{2}}{(T+K)\alpha_{2}^{2}}$ (89)
also, $R_{2}$ can be bounded by a logarithmic dependence of $T$
$R_{2}\leq\sum_{t=0}^{T-1}[\eta_{t}C_{1}\varepsilon_{\text{bias}}]=C_{1}\varepsilon_{\text{bias}}\sum_{t=0}^{T-1}\frac{1}{\alpha_{2}(t+K)}\leq\frac{C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (90)
Combining (89) and (90) with (88) would obtain the following:
$\mathbb{E}\Bigl{[}M_{\overline{E}}(x^{T}-x^{*})\Bigr{]}\leq\frac{K^{2}}{(T+K)^{2}}M_{\overline{E}}(x^{0}-x^{*})+\frac{4\alpha_{2}}{(T+K)\alpha_{2}^{2}}+\frac{C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (91)
Combining (91) with (67) yields (86). ∎
Appendix E Uncertainty Set Support Function Estimators
E.1 Proof of Theorem 3
We have
$\mathbb{E}[M]$ $=\sum_{n=0}^{N_{\max}-1}2^{n+1}\,\mathbb{P}(N^{\prime}=n)+2^{N_{\max}+1}\mathbb{P}(N^{\prime}=N_{\max})$
$=\sum_{n=0}^{N_{\max}-1}2^{n+1}\,\mathbb{P}(N=n)+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=\sum_{n=0}^{N_{\max}-1}\Bigl{(}\frac{2^{n+1}}{2^{n+1}}\Bigr{)}+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=N_{\max}+2^{N_{\max}+1}\mathbb{P}(N\geq N_{\max})$
$=N_{\max}+\frac{2^{N_{\max}+1}}{2^{N_{\max}}}$
$=N_{\max}+2=\operatorname*{\mathcal{O}}(N_{\max}).$ (92)
E.2 Proof of Theorem 4
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untruncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From Wang et al. (2023b), under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
$\mathbb{E}$ $\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]=\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]-\mathbb{E}\left[\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)\right]$
$=\mathbb{E}\left[\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right]-\mathbb{E}\left[\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N+1}}(V)+\frac{\Delta_{N}(V)}{\mathbb{P}(N=n)}\right]$
$=\mathbb{E}\left[\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right]-\mathbb{E}\left[\frac{\Delta_{N}(V)}{\mathbb{P}(N=n)}\right]$
$=\sum_{n=0}^{N_{\max}}\Delta_{n}(V)-\sum_{n=0}^{\infty}\Delta_{n}(V)$
$=\sum_{n=N_{\max}+1}^{\infty}\Delta_{n}(V)$ (93)
For each $\Delta_{n}(V)$, the expectation of absolute value can be bounded as
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]=\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]$
$+\frac{1}{2}\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a,E}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]+\frac{1}{2}\mathbb{E}\left[\left|\sigma_{\hat{\mathsf{P}}^{a,O}_{s,n+1}}(V)-\sigma_{{\mathsf{P}}^{a}_{s}}(V)\right|\right]$ (94)
By the binomial concentration and the Lipschitz property of the support function as in Lemma 2, we know for TV distance uncertainty, we have
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6(1+\frac{1}{\delta})2^{-\frac{n}{2}}\|V\|_{\mathrm{sp}}$ (95)
and for Wasserstein disance uncertainty, we have
$\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6\cdot 2^{-\frac{n}{2}}\|V\|_{\mathrm{sp}}$ (96)
Thus, for TV distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq\sum_{n=N_{\max}+1}^{\infty}\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (97)
and for Wasserstein distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq\sum_{n=N_{\max}+1}^{\infty}\mathbb{E}\left[\left|\Delta_{n}(V)\right|\right]\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}$ (98)
E.3 Proof of Lemma 2
For TV uncertainty sets, for a fixed $V$, for any $p\in\Delta(\mathcal{S})$, define $f_{p}(\mathbf{\mu})\coloneqq p(V-\mathbf{\mu})-\delta\|V-\mathbf{\mu}\|_{\mathrm{sp}}$ and $\mathbf{\mu}_{p}^{*}\coloneqq\arg\max_{\mathbf{\mu}\geq\mathbf{0}}f_{p}(\mathbf{\mu})$. Thus, we have
$\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)=f_{p}(\mu_{p}^{*})-f_{q}(\mu_{q}^{*})$ (99)
since, $\mathbf{\mu}_{p}^{*}$ and $\mathbf{\mu}_{q}^{*}$ are maximizers of $f_{p}$ and $f_{q}$ respectively, we further have
$f_{p}(\mathbf{\mu}_{q}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})\leq f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})\leq f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{p}^{*})$ (100)
Combing (99) and (100) we thus have:
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|$ $\leq\max\{|f_{p}(\mathbf{\mu}_{p}^{*})-f_{q}(\mathbf{\mu}_{p}^{*})|,|f_{p}(\mathbf{\mu}_{q}^{*})-f_{q}(\mathbf{\mu}_{q}^{*})|\}$
$=\max\{|(p-q)(V-\mathbf{\mu}_{p}^{*})|,|(p-q)(V-\mathbf{\mu}_{q}^{*})|\}$ (101)
Note that $\sigma_{\mathcal{P}_{TV}}(V)$ can also be expressed as $\sigma_{\mathcal{P}_{TV}}(V)=p\mathbf{x}^{*}-\delta\|\mathbf{x}^{*}\|_{\mathrm{sp}}$ where $\mathbf{x}^{*}\coloneqq\arg\max_{\mathbf{x}\leq V}(p\mathbf{x}-\delta\|\mathbf{x}\|_{\mathrm{sp}})$. Let $M\coloneqq\max_{s}\mathbf{x}^{*}(s)$ and $m\coloneqq\min_{s}\mathbf{x}^{*}(s)$, then $\|\mathbf{x}\|_{\mathrm{sp}}=M-m$. Denote $\mathbf{e}$ as the all-ones vector, then $\mathbf{x}=\min_{s}V(s)\cdot\mathbf{e}$ is a feasible solution. Thus,
$p\mathbf{x}^{*}-\delta(M-m)\geq p(\min_{s}V(s)\cdot\mathbf{e})-\delta\|\min_{s}V(s)\cdot\mathbf{e}\|_{\mathrm{sp}}=\min_{s}V(s)$ (102)
Since $p$ is a probability vector, $p\mathbf{x}^{*}\leq M$, using the fact that $\delta>0$, we then obtain
$M-\delta(M-m)\geq\min_{s}V(s)\Rightarrow M-m\leq\frac{M-\min_{s}V(s)}{\delta}$ (103)
Since $\mathbf{x^{*}}$ is a feasible solution, we have
$M\leq\max_{s}V(s)\Rightarrow M-\min_{s}V(s)\leq\max_{s}V(s)-\min_{s}V(s)=\|V\|_{\mathrm{sp}}$ (104)
Combining (103) and (104) we obtain
$M-m\leq\frac{\|V\|_{\mathrm{sp}}}{\delta}\Rightarrow m\geq M-\frac{\|V\|_{\mathrm{sp}}}{\delta}\geq\min_{s}V(s)-\frac{\|V\|_{\mathrm{sp}}}{\delta}$ (105)
Where the last inequality is from $M\geq\min_{s}V(s)$, which is a direct result of (103) and the term $\delta(M-m)$ being positive. We finally arrive with
$\mathbf{x}^{*}(j)\in[m,M]\subseteq\Big{[}\min_{s}V(s)-\frac{\|V\|_{\mathrm{sp}}}{\delta},\max_{s}V(s)\Big{]}\quad\forall j\in\mathcal{S}$ (106)
Thus, $\|\mathbf{x}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}$, which leads to
$\|V-\mathbf{\mu}_{p}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}},\quad\|V-\mathbf{\mu}_{q}^{*}\|_{\mathrm{sp}}\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}$ (107)
Combining (107) with (E.3) we obtain the first part of (23).
For Wasserstein uncertainty sets, note that for any $p\in\Delta(\mathcal{S})$ and value function $V$,
$\sigma_{\mathcal{P}_{W}}(V)=\sup_{\lambda\geq 0}\left(\overbrace{-\lambda\delta^{l}+\mathbb{E}_{p}\big{[}\underbrace{\inf_{y\in\mathcal{S}}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}}_{\phi(s,\lambda)}\big{]}}^{g(\lambda,p)}\right).$ (108)
Note that
$\inf_{y\in\mathcal{S}}V(y)\leq\phi(s,\lambda)\leq V(s)+\lambda d(S,s)^{l}=V(s)$ (109)
where the first inequality is because $\lambda d(S,y)^{l}\geq 0$ for any $d$ and $l$. We can then bound $\phi$ by the span of $V$ as
$|\phi(s,\lambda)|\leq\|V\|_{\mathrm{sp}}\quad\forall\lambda\geq 0$ (110)
We then further have that for any $p,q\in\Delta(\mathcal{S})$ and $\lambda\geq 0$,
$|g(\lambda,p)-g(\lambda,q)|\leq\sum_{s\in\mathcal{S}}|p(s)-q(s)||\phi(s,\lambda)|\leq\|p-q\|_{1}\|V\|_{\mathrm{sp}}$ (111)
using (111) and the fact that $|f(\lambda)-g(\lambda)|\leq\epsilon\Rightarrow|\sup_{\lambda}f(\lambda)-\sup_{\lambda}g(\lambda)|\leq\epsilon$, we obtain the second part of (23).
E.4 Proof of Theorem 5
For all $p\in\Delta(\mathcal{S})$, we have $\sigma_{p}(V)\leq\|V\|_{\mathrm{sp}}$, leading to
$\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]+\|V\|^{2}_{\mathrm{sp}}$ (112)
To bound the second moment, note that
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $=\mathbb{E}\left[\left(\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V)+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq\mathbb{E}\left[\left(\|V\|_{\mathrm{sp}}+\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq 2\|V\|_{\mathrm{sp}}^{2}+2\mathbb{E}\left[\left(\frac{\Delta_{N^{\prime}}(V)}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\right]$
$\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\left(\frac{\mathbb{E}[|\Delta_{n}(V)|]}{\mathbb{P}(N^{\prime}=n)}\right)^{2}\mathbb{P}(N^{\prime}=n)$
$=2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{\mathbb{E}[|\Delta_{n}(V)|]^{2}}{\mathbb{P}(N^{\prime}=n)}$ (113)
Under TV distance uncertainty set, by (95), we further have
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{36(1+\frac{1}{\delta})^{2}2^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (114)
Under Wasserstein distance uncertainty set, by (96), we further have
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{362^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (115)
Appendix F Convergence for Robust TD
F.1 Formal Statement of Theorem 6
The first half of Algorithm 2 (line 1 - line 7) can be treated as a special instance of the SA updates in (66) with the bias and variance of the i.i.d. noise term specified in Section 4. To facilitate deriving the bounds of the noise terms, we first analyze the bounds in terms of the $l_{\infty}$ norm, and then translate the bounds in terms of the $\|\cdot\|_{\mathcal{P}}$ semi-norm to obtain the final results.
We start with analyzing the bias and variance of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ for each $t$. Recall the definition of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is as follows:
$\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)=\sum_{a}\pi(a|s)\big{[}r(s,a)-g_{0}+\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\big{]}\quad\forall s\in\mathcal{S}$
Thus, we have for all $s\in\mathcal{S}$,
$\left|\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right|\leq\sum_{a}\pi(a|s)\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|=\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|$ (116)
Which further implies the bias of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is bounded by the bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})$ as follows:
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|$ (117)
Regarding the variance, note that
$\mathbb{E}$ $\left[(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-{\mathbf{T}}_{g_{0}}(V_{t})(s))^{2}\right]$
$=\left(\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)^{2}+\mathrm{Var}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)$
$\leq\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|^{2}+\mathrm{Var}\left(\sum_{a}\pi(a|s)\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right)$
$=\left|\mathbb{E}[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})]-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right|^{2}+\sum_{a}\pi(a|s)^{2}\mathrm{Var}\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{t})\right)$ (118)
To create an upper bound of $\|V\|_{\mathrm{sp}}$ for all possible $V$, define the mixing time of any $p\in\mathcal{P}$ to be
$t^{p}_{\mathrm{mix}}\coloneqq\arg\min_{t\geq 1}\left\{\max_{\mu_{0}}\left\|(\mu_{0}p_{\pi}^{t})^{\top}-\nu^{\top}\right\|_{1}\leq\frac{1}{2}\right\}$ (119)
where $p_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. By Assumption 1, and Lemma 13,and for any value function $V$, we have
$t^{p}_{\mathrm{mix}}<+\infty\quad\text{and}\quad\|V\|_{\mathrm{sp}}\leq 4t^{p}_{\mathrm{mix}}\leq 4t_{\mathrm{mix}}$ (120)
where we define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma 12, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to Lemma 13 and combining the above result with Lemma 13, we obtain that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]\leq\|V_{t}\|^{2}\leq 16t^{2}_{\mathrm{mix}}$ (121)
Regarding TV distance uncertainty sets, using the property of the bias and variance of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ in Theorem 4 and Theorem 5 while combining them with Lemma 13, we have
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (122)
and
$\mathbb{E}$ $\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$ (123)
Similarly, for Wasserstein distance uncertainty sets, we have
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (124)
and
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304t_{\mathrm{mix}}^{2}N_{\max}$ (125)
In order to translate the above bounds from the $l_{\infty}$ norm into the $\|\cdot\|_{\mathcal{P}}$ norm, recall that in line 7 of Algorithm 2, we chose an anchor state $s_{0}$ set $V_{t}(s_{0})=0$ for all $t$ to avoid ambiguity. We thus can draw the following relationship:
Lemma 9.
Let $x\in\mathbb{R}^{S}$ satisfy $x_{i}=0$ for some fixed index $i$. Then
$\|x\|_{\infty}\leq\|x\|_{\mathrm{sp}}=\max_{1\leq j\leq n}x_{j}-\min_{1\leq j\leq n}x_{j}\leq 2\|x\|_{\infty}.$
Moreover, since all semi-norms with the same kernel spaces are equivalent, there are constants $c_{\mathcal{P}},C_{\mathcal{P}}>0$ so that
$c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq\;C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\quad\forall\,x\in\mathbb{R}^{n},$
then
$c_{\mathcal{P}}\|x\|_{\infty}\leq c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq 2C_{\mathcal{P}}\|x\|_{\infty}.$ (126)
Proof.
Since $x_{i}=0$, for every $j$ we have $-\|x\|_{\infty}\leq x_{j}\leq\|x\|_{\infty}$. Hence
$\max_{j}x_{j}\;\leq\;\|x\|_{\infty},\quad\min_{j}x_{j}\;\geq\;-\,\|x\|_{\infty},$
and so
$\|x\|_{\infty}=\max\{\max_{j}x_{j}-0,0-\min_{j}x_{j}\}\leq\|x\|_{\mathrm{sp}}=\max_{j}x_{j}-\min_{j}x_{j}\leq\|x\|_{\infty}-(-\|x\|_{\infty})=2\,\|x\|_{\infty}.$
Since $\|\cdot\|_{\mathrm{sp}}$ and $\|\cdot\|_{\mathcal{P}}$ both have the same kernel of $\{c\mathbf{e}:c\in\mathbb{R}\}$, by the equivalence of semi-norms, it follows that there exists $c_{\mathcal{P}}$ and $C_{\mathcal{P}}$ such that
$c_{\mathcal{P}}\|x\|_{\infty}\leq c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq 2C_{\mathcal{P}}\|x\|_{\infty}$
as claimed. ∎
With the relationship established above, line 1 - line 7 of Algorithm 2 can be formally treated as a special instance of the SA updates in (66) with $B=0$. We now provide the bias and variance of the i.i.d. noise for the different uncertainty sets discussed using Lemma 9 and the estimation bounds in (121)-(F.1). For contamination uncertainty sets, we have
$\varepsilon^{\text{Cont}}_{\text{bias}}=0\quad\text{and}\quad A^{\text{Cont}}=32C_{\mathcal{P}}^{2}t^{2}_{\mathrm{mix}}$ (127)
for TV distance uncertainty sets, we have
$\varepsilon^{\text{TV}}_{\text{bias}}=48C_{\mathcal{P}}(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (128)
and
$A^{\text{TV}}$ $=2C_{\mathcal{P}}^{2}\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (129)
and for Wasserstein distance uncertainty sets, we have
$\varepsilon^{\text{Wass}}_{\text{bias}}=48C_{\mathcal{P}}\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (130)
and
$A^{\text{Wass}}$ $=2C_{\mathcal{P}}^{2}\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (131)
Theorem 10 (Formal version of Theorem 6).
Let $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq 8c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq\rho_{2}L$, if $V_{t}$ is generated by Algorithm 2. Define $V^{*}$ to be the anchored robust value function $V^{*}=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$ for some $c$ such that $V^{*}(s_{0})=0$, then under Assumption 1 and if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$, then for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (132)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}$ $\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}$ (133)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (134)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}$ $\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}$ (135)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (136)
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}},c_{\mathcal{P}}$ are defined in Lemma 9.
Proof.
By Lemma 9 and (120), we have that for any value function $V$ its $\|\cdot\|_{\mathcal{P}}$ norm is bounded as follows:
$\|V\|_{\mathcal{P}}\leq 4C_{\mathcal{P}}t_{\rm mix}$ (137)
Substituting the terms of (127)-(F.1), (137), and Theorem 2 to Theorem 9, we would have for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{8A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (138)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{8A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}$ (139)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (140)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{4A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}$ (141)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{t^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (142)
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}}$ is defined in Lemma 9. We now translate the result back to the standard $l_{\infty}$ norm by applying Lemma 9 again to thee above, we obtain the desired results. ∎
F.2 Proof of Theorem 6
We use the result from Theorem 10, to set $\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 3, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
[…]
F.3 Formal Statement of Theorem 7
To analyze the second part (line 8 - line 14) of Algorithm 2 and provide the provide the complexity for $g_{t}$, we first define the noiseless function $\bar{\delta}(V)$ as
$\bar{\delta}(V)\coloneqq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\big{[}r(s,a)+{\sigma}_{\mathcal{P}^{a}_{s}}(V)\big{]}-V(s)\right)$ (143)
Thus, we have
$\bar{\delta}_{t}=\bar{\delta}(V_{T})+\nu_{t}$ (144)
where $\nu_{t}$ is the noise term with bias equal to the bias $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})$
$\mathbb{E}[\left|\nu_{t}\right|]=\frac{1}{S}\sum_{s}\sum_{a}\left(\pi(a|s)\mathbb{E}\big{[}\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right|\big{]}\right)=\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|$ (145)
By the Bellman equation in Theorem 1, we have $g^{\pi}_{\mathcal{P}}=\bar{\delta}(V^{*})$, which implies
$\left|\bar{\delta}(V_{T})-g^{\pi}_{\mathcal{P}}\right|$ $=\left|\bar{\delta}(V_{T})-\bar{\delta}(V^{*})\right|$
$\leq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V^{*})\right|+\left|V_{T}(s)-V^{*}(s)\right|\right)$
$\leq\frac{1}{S}\sum_{s}\left(\sum_{a}\pi(a|s)\left|{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V^{*})\right|+\left|V_{T}(s)-V^{*}(s)\right|\right)$
$\leq\frac{1}{S}\sum_{s}2\|V_{T}-V^{*}\|_{\mathrm{sp}}$
$=2\|V_{T}-V^{*}\|_{\mathrm{sp}}$
$\leq 4\|V_{T}-V^{*}\|_{\infty}$ (146)
Where the last inequality is by Lemma 9. Thus, the following recursion can be formed
$\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|$ $=\left|g_{t}+\beta_{t}(\bar{\delta}_{t}-g_{t})-g^{\pi}_{\mathcal{P}}\right|$
$=\left|g_{t}-g^{\pi}_{\mathcal{P}}+\beta_{t}(\bar{\delta}_{t}-g^{\pi}_{\mathcal{P}}+g^{\pi}_{\mathcal{P}}-g_{t})\right|$
$=\left|g_{t}-g^{\pi}_{\mathcal{P}}+\beta_{t}(\bar{\delta}(V^{T})-g^{\pi}_{\mathcal{P}}+\nu_{t}+g^{\pi}_{\mathcal{P}}-g_{t})\right|$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(\left|\bar{\delta}(V^{T})-g^{\pi}_{\mathcal{P}}\right|+\left|\nu_{t}\right|)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(4\|V_{T}-V^{*}\|_{\infty}+\left|\nu_{t}\right|)$ (147)
Thus, taking expectation conditioned on the filtration $\mathcal{F}^{t}$ yields
$\mathbb{E}$ $\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (148)
By letting $\zeta_{t}\coloneqq\Pi_{i=0}^{t-1}(1-\beta_{t})$, we obtain the $T$-step recursion as follows:
$\mathbb{E}$ $\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|$
$+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (149)
By setting $\beta_{t}\coloneqq\frac{1}{t+1}$, we have $\zeta_{T}=\frac{1}{T+1}\leq\frac{1}{T}$ and $\sum_{t=0}^{T-1}\beta_{t}\leq 2\log T$, (149) implies
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(8\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (150)
Theorem 11 (Formal version of Theorem 7).
Following all notations and assumptions in Theorem 10, then for contamination uncertainty sets,
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right)$ (151)
for TV distance uncertainty sets,
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (152)
for Wasserstein distance uncertainty sets,
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (153)
Where all variables are defined the same as in Theorem 10.
Proof.
By Theorem 10, taking square root on both side and utilizing the concavity of square root function, we have for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}$ (154)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}}$ (155)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}}$ (156)
Regarding the bound for the absolute bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}$, from Lemma 12, we have for contamination uncertainty,
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|=0$ (157)
In addition, combining (97)-(98) with Lemma 13, we have for for TV distance uncertainty,
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (158)
and for Wasserstein distance uncertainty, we have
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (159)
Combining (154)-(159) with (150) gives the desired result. ∎
F.4 Proof of Theorem 7
We use the result from Theorem 11, to set $\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 3, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
Appendix G Some Auxiliary Lemmas for the Proofs
Lemma 10 (Theorem IV in Berger and Wang (1992)).
Let $\mathcal{Q}$ be a bounded set of square matrix such that $\rho(Q)<\infty$ for all $Q\in\mathcal{Q}$ where $\rho(\cdot)$ denotes the spectral radius. Then we have
$\sup_{Q\in\mathcal{Q}}\rho(Q)=\lim_{k\rightarrow\infty}\sup_{Q_{i}\in\mathcal{Q}}\rho(Q_{k}\ldots Q_{1})^{\frac{1}{k}}$ (160)
Lemma 11 (Lemma 6 in Zhang et al. (2021)).
Under the setup and notation in Appendix D.1.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$, we have
$\mathbb{E}\left[\|x^{t+1}-x^{t}\|^{2}_{s,\overline{E}}|\mathcal{F}^{t}\right]\leq(16+4B)c_{u}\rho_{2}\eta_{t}^{2}M_{\overline{E}}(x^{t}-x^{*})+2A\rho_{2}\eta_{t}^{2}.$ (161)
Lemma 12 (Theorem D.1 in Wang et al. (2023b)).
The estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (18) for contamination uncertainty sets is unbiased and has bounded variance as follows:
$\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]={\sigma}_{\mathcal{P}^{a}_{s}}(V),\quad\text{and}\quad\mathrm{Var}(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V))\leq\|V\|^{2}$ (162)
Lemma 13 (Ergodic case of Lemma 9 in Wang et al. (2022)).
For any average reward MDP with stationary policy $\pi$ and the mixing time defined as
$\tau_{\mathrm{mix}}\coloneqq\arg\min_{t\geq 1}\left\{\max_{\mu_{0}}\left\|(\mu_{0}P_{\pi}^{t})^{\top}-\nu^{\top}\right\|_{1}\leq\frac{1}{2}\right\}$ (163)
where $P_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. If $P_{\pi}$ is irreducible and aperiodic, then $\tau_{\mathrm{mix}}<+\infty$ and for any value function $V$, we have
$\|V\|_{\mathrm{sp}}\leq 4\tau_{\mathrm{mix}}$ (164)